### v2

# Heat Flow under Semi-Flat Collapse with Conic Renormalization

Abstract.
Motivated by the Strominger–Yau–Zaslow (SYZ) collapse of elliptic K3 surfaces, we study heat flow on Ricci-flat Kähler manifolds $(X_{t},g(t))$ equipped over $B_{\mathrm{reg}}=B\setminus D$ with a semi-flat $2$-torus fibration $\Pi_{t}:X_{t}\to B$. Assuming an exponentially accurate semi-flat product approximation and a uniform vertical spectral gap at the collapse scale, we prove that for each fixed $\tau>0$ the fiber-compressed heat operators $P_{t}e^{-\tau\Delta_{g(t)}}I_{t}$ converge strongly on $L^{2}$ to the base heat semigroup $e^{-\tau\Delta_{g_{B}}}$; equivalently, bilinear pairings of the total-space heat kernel against fiber-constant test functions converge on compact subsets of $B_{\mathrm{reg}}$.
Near the discriminant $D$ the limiting base metric is conic. We introduce a canonical renormalized bilinear functional $K_{B}^{\mathrm{ren}}$ by replacing the pairing in neighborhoods of each conic point with the model flat-cone heat kernel. A conic parametrix proves cutoff- and chart-independence of $K_{B}^{\mathrm{ren}}$ and shows that the total-space pairings converge to $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ for all $\Phi,\Psi\in C_{c}^{\infty}(B)$. Under exponentially small geometric errors we obtain an explicit exponential convergence rate, uniform for $\tau$ in compact subsets of $(0,\infty)$, and we also treat Neumann and nonnegative Robin boundary conditions.
1. Introduction and statement of results
1.1. Motivation: heat flow under SYZ-type collapse
Collapsing families of Ricci–flat Kähler metrics arise naturally in complex geometry and mirror symmetry, most prominently in the Strominger–Yau–Zaslow (SYZ) picture [1] and in the degeneration theory of elliptically fibered K3 surfaces. In such settings one expects that, away from singular fibers, the geometry becomes increasingly close to a semi-flat product: a fixed base metric coupled with flat torus fibers whose diameters shrink to zero, with “instanton corrections” that are exponentially small in the collapse scale; see, for instance, [2, 10, 11].Analytically, this collapse suggests a clean separation between low-energy horizontal dynamics and high-energy vertical oscillations, governed by a uniform fiberwise spectral gap. Such spectral separation is consistent with the adiabatic limit framework developed in [10, 11].
The purpose of this paper is to make this picture precise at the level of heat flow. Since pointwise convergence of heat kernels is generally unstable under collapse, we focus on a robust notion: bilinear heat-kernel pairings tested against fiber-constant lifts of base functions, or equivalently the strong convergence of fiberwise-compressed heat semigroups. The new feature occurs at the discriminant: the limiting base geometry is conic near $D$, and to treat test functions that meet $D$ one needs an intrinsic way to pass from purely interior convergence on $B_{\mathrm{reg}}$ to a global, cutoff-independent limiting bilinear object. We introduce such a limit functional by a canonical conic renormalization based on the flat-cone model.
1.2. Geometric setting and analytic identifications
Let $B$ be a smooth surface, let $D=\{p_{1},\dots,p_{N}\}\subset B$ be a finite discriminant set, and set
$B_{\mathrm{reg}}:=B\setminus D.$
Let $(X_{t},g(t))$ be a family of Riemannian manifolds equipped with smooth surjective submersions $\Pi_{t}:X_{t}\to B$ whose fibers $F_{b,t}:=\Pi_{t}^{-1}(b)$ are flat $2$-tori over $B_{\mathrm{reg}}$. Fix a precompact domain $B_{0}\Subset B_{\mathrm{reg}}$ with smooth boundary and set $X_{t,0}:=\Pi_{t}^{-1}(B_{0})$. We impose a boundary condition (Dirichlet or Neumann; Robin with nonnegative weight is discussed later) on $\partial B_{0}$ and simultaneously on $\partial X_{t,0}=\Pi_{t}^{-1}(\partial B_{0})$.
[…]
Figure 1. Schematic picture of the fibration $\Pi_{t}:X_{t}\to B$ with finite discriminant set $D$. Over the regular locus $B_{\mathrm{reg}}=B\setminus D$, the fibers are flat $2$-tori $\mathbb{T}^{2}$ that collapse as $t\downarrow 0$.
Semi-flat control and vertical gap (interior).
On each $B_{0}\Subset B_{\mathrm{reg}}$ we assume a semi-flat product approximation
$g(t)=g_{B}\oplus g_{F,t}+E_{t}\quad\text{on }X_{t,0},$
where $g_{B}$ is a fixed smooth metric on $B_{0}$, $g_{F,t}|_{F_{b,t}}$ is the flat metric on the fiber, and the error $E_{t}$ is exponentially small in a collapse scale $s^{*}(t)\downarrow 0$:
$\|E_{t}\|_{C^{1}(X_{t,0})}\leq\varepsilon_{t},\qquad\varepsilon_{t}\leq Ce^{-c/s^{*}(t)}.$
We also assume a uniform vertical spectral gap at the collapse scale:
$\lambda_{1}(F_{b,t},g_{F,t})\ \geq\ c_{0}\,s^{*}(t)^{-2},\qquad\forall b\in B_{0},$
where $\lambda_{1}$ denotes the first nonzero eigenvalue of the fiber Laplacian. These hypotheses are standard in semi-flat regimes arising in SYZ-type degenerations [2, 3, 4].
Normalized lift/average.
A key point is to compare operators on the varying spaces $X_{t,0}$ with operators on the fixed base $B_{0}$. Using quantitative fiberwise disintegration of the volume (available under the semi-flat control), we define the normalized lift $I_{t}:L^{2}(B_{0})\to L^{2}(X_{t,0})$ and fiber-average $P_{t}:L^{2}(X_{t,0})\to L^{2}(B_{0})$ so that
$P_{t}I_{t}=\mathrm{Id}_{L^{2}(B_{0})},\qquad I_{t}=P_{t}^{*},$
and $I_{t},P_{t}$ are near-isometries on $L^{2}$ (and compatible with horizontal $H^{1}$) with errors $O(\varepsilon_{t})$. Heuristically, $I_{t}$ identifies a base function with a fiber-constant function, normalized by the fiber area.
1.3. Bilinear pairings and compressed semigroups
Let $H_{t,0}:=-\Delta_{g(t)}$ on $L^{2}(X_{t,0})$ and $H_{B,0}:=-\Delta_{g_{B}}$ on $L^{2}(B_{0})$, with the chosen boundary condition. Denote the heat semigroups by
$T_{t}(\tau):=e^{-\tau H_{t,0}},\qquad T_{B}(\tau):=e^{-\tau H_{B,0}},$
with corresponding heat kernels $K_{t}(\cdot,\cdot;\tau)$ and $K_{B}(\cdot,\cdot;\tau)$. For $\Phi,\Psi\in C_{c}^{\infty}(B_{0})$, let $\Phi_{t}:=I_{t}\Phi$ and $\Psi_{t}:=I_{t}\Psi$. The bilinear pairing of $K_{t}$ against fiber-constant lifts can be written equivalently as
(1) $\iint_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,\Phi_{t}(x)\Psi_{t}(y)\,d\mu_{g(t)}(x)d\mu_{g(t)}(y)\ =\ \langle\Phi,\ (P_{t}T_{t}(\tau)I_{t})\Psi\rangle_{L^{2}(B_{0})}.$
Thus strong convergence of the compressed semigroups $P_{t}T_{t}(\tau)I_{t}$ implies convergence of bilinear heat-kernel pairings.
1.4. Discriminant regime and intrinsic conic renormalization
The previous discussion applies on $B_{\mathrm{reg}}$. To treat test functions that may meet the discriminant, we incorporate the conic nature of the limiting base geometry near $D$.
Conic model on the base.
For each $p_{j}\in D$, choose a wedge chart $U_{j}\ni p_{j}$ with polar coordinates $(r,\theta)$ centered at $p_{j}$, where $r=\mathrm{dist}_{g_{B}}(\cdot,p_{j})$. We assume $g_{B}$ is $C^{1}$-close to the flat cone
$g_{j}^{\mathrm{cone}}:=dr^{2}+\alpha_{j}^{2}r^{2}d\theta^{2}\quad\text{on }U_{j}\setminus\{p_{j}\},$
in a quantitative sense (cf. Assumption 5.1), for some cone factors $\alpha_{j}>0$. Let $K^{\mathrm{cone}}_{j}(\cdot,\cdot;\tau)$ denote the Friedrichs heat kernel on $(U_{j},g_{j}^{\mathrm{cone}})$, which admits an explicit Bessel expansion and sharp Gaussian-type bounds [5].
Why renormalization is necessary.
When $\Phi,\Psi$ do not vanish at $D$, the interior convergence on precompact subsets of $B_{\mathrm{reg}}$ does not by itself produce an intrinsic global limit: truncating away a neighborhood of $D$ and sending the cutoff radius to zero creates a local contribution near each conic point that must be accounted for in a canonical way. The flat-cone model provides the correct local replacement, and leads to a cutoff-independent target functional.
Renormalized bilinear functional.
Fix a smooth radial cutoff $\eta_{\rho}$ supported near $r\sim\rho$ and write $\Phi=\Phi^{(\rho)}+\Phi^{<\rho}$ with $\Phi^{(\rho)}:=\eta_{\rho}\Phi$ and $\mathrm{supp}(\Phi^{<\rho})\subset\{r<\rho\}$, and similarly for $\Psi$. Choose cutoffs $\chi_{j}\in C_{c}^{\infty}(U_{j})$ with $\chi_{j}\equiv 1$ near $p_{j}$. We define
(2) $\begin{split}K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau):=\lim_{\rho\downarrow 0}\Bigg[&\iint_{B\times B}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\Psi^{(\rho)}(b^{\prime})\\ &\qquad d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})\\ &+\sum_{j=1}^{N}\iint_{U_{j}\times U_{j}}(\chi_{j}\otimes\chi_{j})\,K^{\mathrm{cone}}_{j}\\ &\qquad\Phi^{<\rho}\otimes\Psi^{<\rho}\,d\mu^{\otimes 2}_{g^{\mathrm{cone}}_{j}}\Bigg].\end{split}$
A standard conic parametrix for $K_{B}$ near $p_{j}$ implies that the limit exists and is independent of the choices of $\eta_{\rho}$ and $\chi_{j}$; moreover if $\mathrm{supp}\,\Phi\cup\mathrm{supp}\,\Psi\subset B_{\mathrm{reg}}$, then for $\rho$ small one has $\Phi^{<\rho}=\Psi^{<\rho}=0$ and the renormalization is trivial, so that $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ reduces to the usual base heat-kernel pairing.
1.5. Main result: renormalized bilinear limit across the discriminant
We can now state the main theorem in a form suited for applications.
Theorem 1.1 (Renormalized bilinear heat-kernel limit across the discriminant).
Assume the semi-flat control and the vertical spectral gap on every precompact subset of $B_{\mathrm{reg}}$, and assume that near each $p_{j}\in D$ the base metric $g_{B}$ is $C^{1}$-close to the flat cone $g_{j}^{\mathrm{cone}}$ with cone factor $\alpha_{j}$. Let $\Phi,\Psi\in C_{c}^{\infty}(B)$ and $\tau>0$. Then, with the normalized lifts $\Phi_{t}:=I_{t}\Phi$, $\Psi_{t}:=I_{t}\Psi$,
(3) $\lim_{t\downarrow 0}\ \iint_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,\Phi_{t}(x)\Psi_{t}(y)\,d\mu_{g(t)}(x)d\mu_{g(t)}(y)\ =\ K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau),$
where $K_{B}^{\mathrm{ren}}$ is the canonical renormalized functional defined in (2). If $\mathrm{supp}\,\Phi\cup\mathrm{supp}\,\Psi\subset B_{\mathrm{reg}}$, then $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ coincides with the usual base pairing and (3) reduces to the interior bilinear limit.
If, in addition, the semi-flat errors are exponentially small (e.g. $\|E_{t}\|_{C^{1}}\leq Ce^{-c/s^{*}(t)}$), then the convergence in (3) holds with an explicit exponential rate, uniform for $\tau$ in compact subintervals of $(0,\infty)$.
Remark 1.2 (Compressed semigroups and bilinear kernels).
On precompact $B_{0}\Subset B_{\mathrm{reg}}$, the bilinear limit is equivalent to strong convergence of the compressed heat semigroups
$P_{t}e^{-\tau\Delta_{g(t)}}I_{t}\ \longrightarrow\ e^{-\tau\Delta_{g_{B}}}\quad\text{on }L^{2}(B_{0}),$
and can be derived via Mosco convergence of the associated Dirichlet forms. In the discriminant regime, the renormalized functional (2) provides a canonical way to combine interior limits on $B_{\mathrm{reg}}$ with the flat-cone model near $D$; in particular the limiting bilinear object is independent of the auxiliary cutoffs used to localize near the conic points.
1.6. Idea of the proof
The argument separates naturally into an interior collapse analysis and a conic discriminant analysis.
(i) Interior collapse and fiberwise compression. On each $B_{0}\Subset B_{\mathrm{reg}}$, the semi-flat control provides quantitative disintegration of $d\mu_{g(t)}$ along the fibers and near-isometric identifications $(I_{t},P_{t})$. The vertical spectral gap yields a Poincaré inequality that suppresses fiber-mean-zero components and leads to a coercive splitting of the Dirichlet energy into a fiber-constant part and a vertical part. This gives Mosco convergence of Dirichlet forms under the identifications, hence strong resolvent/semigroup convergence by standard form methods [14, 13, 15]. The adjoint relation $I_{t}=P_{t}^{*}$ and kernel representations then convert operator limits into bilinear kernel limits as in (1), with dominated convergence justified by uniform Gaussian bounds on bounded-geometry domains [7, 8, 21].
(ii) Exhaustion and boundary decoupling. To globalize from $B_{0}\Subset B_{\mathrm{reg}}$ to $B_{\mathrm{reg}}$, we use a smooth exhaustion and Davies–Gaffney off-diagonal estimates [7, 16] to show that artificial boundary effects are exponentially small in the exhaustion index on fixed time windows.
(iii) Conic parametrix and intrinsic renormalization. Near each $p_{j}\in D$, we compare the base heat kernel to the flat-cone heat kernel via a conic parametrix [9, 22, 5]. The remainder has an integrable defect controlled by the $C^{1}$-closeness to the cone. This yields existence and cutoff-independence of $K_{B}^{\mathrm{ren}}$ and, combined with the interior convergence on $\{r\geq\rho\}$, produces the renormalized limit (3) after sending first $t\downarrow 0$ and then $\rho\downarrow 0$. Quantitative bookkeeping is obtained by combining the interior convergence rate, off-diagonal (Davies–Gaffney) decoupling across a $\rho$-buffer, and the conic parametrix remainder.
1.7. Further consequences and organization of the paper
The renormalized bilinear limit has a number of standard spectral consequences after compression, such as stability of localized spectral data and convergence of localized heat traces. In order to keep the present paper focused on the discriminant mechanism and the intrinsic renormalization, we confine ourselves to a minimal set of such corollaries.
Organization. Section 2 develops the semi-flat disintegration and the normalized identifications. Section 3 proves Mosco convergence of the relevant Dirichlet forms on $B_{0}\Subset B_{\mathrm{reg}}$. Section 4 derives strong convergence of compressed heat semigroups and the interior bilinear kernel limit. Section 5 constructs a conic parametrix on the base, defines the cutoff-independent functional $K_{B}^{\mathrm{ren}}$, and identifies it as the global limit object across the discriminant. Section 6 records quantitative error estimates and an iterated-limit statement. Finally, Section 8 summarizes verifiable geometric hypotheses in semi-flat elliptic fibrations that imply the analytic assumptions used throughout.
2. Semi-flat collapse and normalized identifications
This section fixes notation and records the quantitative identifications that allow us to compare analysis on the varying total spaces $(X_{t},g(t))$ with analysis on the fixed base $(B,g_{B})$. The key output is a canonical $L^{2}$ identification pair
$I_{t}:L^{2}(B_{0})\to L^{2}(X_{t,0}),\qquad P_{t}:L^{2}(X_{t,0})\to L^{2}(B_{0}),$
satisfying $P_{t}I_{t}=\mathrm{Id}$ and $I_{t}=P_{t}^{*}$, where $I_{t}$ lifts base functions to fiber-constant functions and $P_{t}$ averages along fibers with the correct normalization. In addition, $I_{t}$ is compatible with the horizontal $H^{1}$ energy up to $O(\varepsilon_{t})$ errors under the semi-flat $C^{1}$ control.
2.1. Geometric setting on the regular locus
Let $(B,g_{B})$ be a smooth Riemannian surface and let $D=\{p_{1},\dots,p_{N}\}\subset B$ be a finite set. We set $B_{\mathrm{reg}}:=B\setminus D$. For each $t\in(0,t_{0}]$ we are given a smooth surjective submersion
$\Pi_{t}:X_{t}\longrightarrow B_{\mathrm{reg}}$
whose fibers $F_{b,t}:=\Pi_{t}^{-1}(b)$ are flat $2$-tori.
Fix once and for all a precompact domain $B_{0}\Subset B_{\mathrm{reg}}$ with smooth boundary. We denote the restricted total space by
$X_{t,0}:=\Pi_{t}^{-1}(B_{0})\Subset X_{t}.$
Throughout the paper we impose the same boundary type on $\partial B_{0}$ and on $\partial X_{t,0}=\Pi_{t}^{-1}(\partial B_{0})$; for definiteness this section is written for Dirichlet boundary conditions (the Neumann and nonnegatively weighted Robin cases are addressed later, and the constructions below remain unchanged).
2.2. Semi-flat product control and collapse scale
We assume that on $X_{t,0}$ the metric $g(t)$ is close to a product metric with exponentially small error. More precisely, we fix:
Assumption 2.1 (Semi-flat $C^{1}$ product control).
There exist the following objects: (i) a fixed smooth base metric $g_{B}$ on $B_{0}$; (ii) a family of flat metrics $g_{F,t}$ along the fibers $F_{b,t}$; (iii) a product reference metric $g_{t}^{\Pi}:=g_{B}\oplus g_{F,t}$ on $X_{t,0}$ (with respect to a horizontal/vertical splitting); (iv) a collapse scale $s(t)\downarrow 0$ as $t\downarrow 0$; and (v) an error gauge $\varepsilon_{t}\geq 0$ with $\varepsilon_{t}\to 0$. such that on $X_{t,0}$ we have
(4) $g(t)=g_{t}^{\Pi}+E_{t},\qquad\|E_{t}\|_{C^{1}(X_{t,0},\,g_{t}^{\Pi})}\leq\varepsilon_{t},$
and the fiber diameters satisfy
(5) $\mathrm{diam}_{g_{F,t}}(F_{b,t})\leq s(t)\qquad\text{for all }b\in B_{0}.$
Such semi-flat/adiabatic product regimes arise in geometric collapse settings (e.g. elliptic K3 and abelian-fibred Calabi–Yau cases); see [2, 10, 11].
Assumption 2.2 (Uniform vertical spectral gap).
There exists $c_{\mathrm{gap}}>0$ such that for all $b\in B_{0}$ and all $t\in(0,t_{0}]$,
(6) $\lambda_{1}(F_{b,t},g_{F,t})\ \geq\ c_{\mathrm{gap}}\,s(t)^{-2},$
where $\lambda_{1}$ is the first nonzero eigenvalue of the (nonnegative) Laplacian on the flat torus fiber.
Assumption 2.3 (Bounded geometry on the base).
The domain $(B_{0},g_{B})$ has positive injectivity radius and uniformly bounded curvature (with finitely many covariant derivatives, sufficient for standard heat kernel bounds on $B_{0}$).
Remark 2.4 (On rates).
In applications to SYZ-type collapse one often has $\varepsilon_{t}\leq Ce^{-c/s(t)}$. In the core convergence results we only use $\varepsilon_{t}\to 0$; the exponential bound is only needed when we state explicit rates.
2.3. Quantitative disintegration of volume
Write $\mu_{t}:=\mu_{g(t)}$ for the Riemannian measure on $(X_{t,0},g(t))$ and $\mu_{B}:=\mu_{g_{B}}$ for the base measure on $(B_{0},g_{B})$. For each fiber $(F_{b,t},g_{F,t})$ we denote by $d\vartheta_{b,t}$ the Haar probability measure. By disintegration of $\mu_{t}$ along the submersion $\Pi_{t}$ (coarea) [12], there exist measurable functions $A_{t}:B_{0}\to(0,\infty)$ and $\rho_{t}:X_{t,0}\to(0,\infty)$ such that for all $f\in L^{1}(X_{t,0},\mu_{t})$,
(7) $\int_{X_{t,0}}f\,d\mu_{t}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b),$
and $\rho_{t}$ is normalized fiberwise:
(8) $\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1\qquad\text{for all }b\in B_{0}.$
Heuristically, $A_{t}(b)$ is the fiber area (at base point $b$) and $\rho_{t}$ measures the deviation of $\mu_{t}$ from the product disintegration relative to $g_{B}\oplus g_{F,t}$.
The semi-flat $C^{1}$ control implies that $\rho_{t}$ is close to $1$ and that $A_{t}$ varies slowly along the base.
Lemma 2.5 (Quantitative control of $(A_{t},\rho_{t})$).
Under Assumption 2.1, there exists $C>0$ (independent of $t$) such that on $X_{t,0}$
(9) $\|\rho_{t}-1\|_{L^{\infty}(X_{t,0})}\ \leq\ C\,\varepsilon_{t},$
and on $B_{0}$
(10) $\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}\ \leq\ C\,\varepsilon_{t}.$
Proof (sketch, with explicit dependencies).
In a local trivialization $X_{t,0}\simeq B_{0}\times\mathbb{T}^{2}$ compatible with the horizontal/vertical splitting, the product reference metric $g_{t}^{\Pi}=g_{B}\oplus g_{F,t}$ induces the exact disintegration
$d\mu_{g_{t}^{\Pi}}=A_{t}^{\Pi}(b)\,d\mu_{B}(b)\,d\vartheta_{b,t},$
with $\rho_{t}\equiv 1$ for the product model. The $C^{1}$-smallness (4) implies that the Radon–Nikodym derivative of $\mu_{t}$ relative to $\mu_{g_{t}^{\Pi}}$ is $1+O(\varepsilon_{t})$ and that its horizontal derivatives are $O(\varepsilon_{t})$. Translating this into the normalized disintegration (7) gives $\rho_{t}=1+O(\varepsilon_{t})$ and $\nabla_{g_{B}}\log A_{t}=O(\varepsilon_{t})$ uniformly on $B_{0}$. A fully detailed coordinate computation is standard in semi-flat settings and is omitted. ∎
2.4. Normalized lift and fiber-average operators
We can now define the canonical identifications between base and total space.
Definition 2.6 (Normalized lift and average).
For $v\in L^{2}(B_{0},\mu_{B})$ define the normalized lift $I_{t}v\in L^{2}(X_{t,0},\mu_{t})$ by
(11) $(I_{t}v)(x):=A_{t}(\Pi_{t}x)^{-1/2}\,v(\Pi_{t}x),$
which is fiber-constant by construction. For $u\in L^{2}(X_{t,0},\mu_{t})$ define the normalized fiber-average $P_{t}u\in L^{2}(B_{0},\mu_{B})$ by
(12) $(P_{t}u)(b):=A_{t}(b)^{1/2}\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}.$
Lemma 2.7 (Algebraic identities and adjointness).
The operators $(I_{t},P_{t})$ satisfy
(13) $P_{t}I_{t}=\mathrm{Id}_{L^{2}(B_{0})},$
and $I_{t}=P_{t}^{*}$ as operators between $L^{2}(B_{0},\mu_{B})$ and $L^{2}(X_{t,0},\mu_{t})$.
Proof.
Let $v\in L^{2}(B_{0})$. By (12), (11), and the normalization (8),
$(P_{t}I_{t}v)(b)=A_{t}(b)^{1/2}\int_{F_{b,t}}A_{t}(b)^{-1/2}v(b)\,\rho_{t}\,d\vartheta_{b,t}=v(b).$
For adjointness, let $\phi\in L^{2}(B_{0})$ and $u\in L^{2}(X_{t,0})$. Using (7), we compute
$\langle I_{t}\phi,u\rangle_{L^{2}(X_{t,0})}$ $=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}A_{t}(b)^{-1/2}\phi(b)\,u\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b)$
$=\int_{B_{0}}\phi(b)\,(P_{t}u)(b)\,d\mu_{B}(b)$
(14) $=\langle\phi,P_{t}u\rangle_{L^{2}(B_{0})}.$
Thus $I_{t}=P_{t}^{*}$. ∎
Lemma 2.8 ($L^{2}$ isometry and contraction).
For all $v\in L^{2}(B_{0})$ and all $u\in L^{2}(X_{t,0})$,
(15) $\|I_{t}v\|_{L^{2}(X_{t,0})}=\|v\|_{L^{2}(B_{0})},$
and
(16) $\|P_{t}u\|_{L^{2}(B_{0})}\leq\|u\|_{L^{2}(X_{t,0})}.$
Proof.
By (11), (7), and (8),
$\|I_{t}v\|_{L^{2}(X_{t,0})}^{2}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}A_{t}(b)^{-1}|v(b)|^{2}\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b)=\int_{B_{0}}|v(b)|^{2}\,d\mu_{B}(b).$
For (16), note that $\rho_{t}\,d\vartheta_{b,t}$ is a probability measure on each fiber, so by Jensen
$|(P_{t}u)(b)|^{2}=A_{t}(b)\left|\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}\right|^{2}\leq A_{t}(b)\int_{F_{b,t}}|u|^{2}\,\rho_{t}\,d\vartheta_{b,t}.$
Integrating in $b$ and using (7) gives $\|P_{t}u\|_{L^{2}(B_{0})}^{2}\leq\|u\|_{L^{2}(X_{t,0})}^{2}$. ∎
2.5. Horizontal $H^{1}$ compatibility for fiber-constant lifts
The collapsed limit is governed by the horizontal energy. Since $I_{t}v$ is fiber-constant, its vertical gradient vanishes. The next lemma quantifies that its energy is close to the base energy, up to a controlled lower-order term coming from the normalization $A_{t}^{-1/2}$.
Lemma 2.9 (Horizontal energy of fiber-constant lifts).
Assume Lemma 2.5 and the $C^{1}$ semi-flat control (4). Then there exists $C>0$ independent of $t$ such that for all $v\in H^{1}(B_{0})$,
(17) $\int_{X_{t,0}}|\nabla I_{t}v|_{g(t)}^{2}\,d\mu_{t}=\int_{B_{0}}|\nabla v|_{g_{B}}^{2}\,d\mu_{B}\ +\ O(\varepsilon_{t})\Big(\|v\|_{H^{1}(B_{0})}^{2}\Big).$
In particular, $I_{t}:H^{1}(B_{0})\to H^{1}(X_{t,0})$ is bounded uniformly in $t$.
Proof (outline with explicit computation).
Work in semi-flat coordinates for $g_{t}^{\Pi}=g_{B}\oplus g_{F,t}$. Since $I_{t}v=A_{t}^{-1/2}\,v\circ\Pi_{t}$, the vertical gradient vanishes for the product metric: $\nabla_{F}(I_{t}v)\equiv 0$. For the horizontal gradient,
$\nabla_{B}(I_{t}v)=A_{t}^{-1/2}\,\nabla v\circ\Pi_{t}\ -\ \frac{1}{2}A_{t}^{-1/2}\,v\circ\Pi_{t}\,\nabla_{B}\log A_{t}.$
Squaring, integrating with $d\mu_{t}$, and using (7) together with $\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1$ yields the identity
$\int_{X_{t,0}}|\nabla_{B}(I_{t}v)|_{g_{B}}^{2}\,d\mu_{t}=\int_{B_{0}}\Big(|\nabla v|_{g_{B}}^{2}-\langle\nabla v,v\,\nabla_{B}\log A_{t}\rangle_{g_{B}}+\frac{1}{4}|v|^{2}|\nabla_{B}\log A_{t}|_{g_{B}}^{2}\Big)\,d\mu_{B}.$
The cross term is bounded by Cauchy–Schwarz and (10), giving an $O(\varepsilon_{t})\|v\|_{H^{1}}^{2}$ contribution. Finally, replacing $g(t)$ by $g_{t}^{\Pi}$ introduces an additional $O(\varepsilon_{t})\|v\|_{H^{1}}^{2}$ error by (4). Combining these estimates gives (17). ∎
2.6. Fiber-constant projection (notation)
We will repeatedly separate fiber-constant and fiber-oscillatory components in $L^{2}(X_{t,0})$. To avoid confusion with the fibration map $\Pi_{t}:X_{t}\to B_{\mathrm{reg}}$, we denote the $L^{2}$ projection onto the fiber-constant sector by $\Pi_{t}^{\mathrm{fc}}$.
Definition 2.10 (Fiber-constant projection).
Define the projection onto the fiber-constant sector by
$\Pi_{t}^{\mathrm{fc}}:=I_{t}P_{t}:L^{2}(X_{t,0})\to L^{2}(X_{t,0}),$
and its complement by $(\Pi_{t}^{\mathrm{fc}})^{\perp}:=\mathrm{Id}-\Pi_{t}^{\mathrm{fc}}$.
By Lemma 2.7, $\Pi_{t}^{\mathrm{fc}}$ is self-adjoint and $(\Pi_{t}^{\mathrm{fc}})^{2}=\Pi_{t}^{\mathrm{fc}}$ on $L^{2}(X_{t,0})$.
2.7. Section summary and next steps
Assumptions 2.12.3 yield a canonical $L^{2}$ identification pair $(I_{t},P_{t})$ with $P_{t}I_{t}=\mathrm{Id}$ and $I_{t}=P_{t}^{*}$, together with $L^{2}$ isometry/contraction (Lemma 2.8) and horizontal $H^{1}$ compatibility for fiber-constant lifts (Lemma 2.9).
In the next section we introduce the Dirichlet forms and generators, use the vertical spectral gap Assumption 2.2 to obtain a vertical Poincaré inequality for $(\Pi_{t}^{\mathrm{fc}})^{\perp}$, and then prove Mosco convergence of Dirichlet forms on the identified spaces. This will lead to strong convergence of the compressed heat semigroups and the interior bilinear heat-kernel limit on $B_{0}\Subset B_{\mathrm{reg}}$.
3. Dirichlet forms, vertical Poincaré inequality, and Mosco convergence
Throughout this section we fix a precompact domain $B_{0}\Subset B_{\mathrm{reg}}$ with smooth boundary and set $X_{t,0}:=\Pi_{t}^{-1}(B_{0})$. We work under Dirichlet boundary conditions on $\partial B_{0}$ and $\partial X_{t,0}$ (the Neumann and nonnegatively weighted Robin cases are treated later and do not affect the arguments below).
3.1. Dirichlet forms on the base and on the total space
Let $\mu_{B}:=\mu_{g_{B}}$ and $\mu_{t}:=\mu_{g(t)}$ denote the Riemannian measures on $B_{0}$ and $X_{t,0}$. We write
$L^{2}_{B}:=L^{2}(B_{0},\mu_{B}),\qquad L^{2}_{t}:=L^{2}(X_{t,0},\mu_{t}).$
Base Dirichlet form.
Define the closed Dirichlet form
(18) $E_{B}[v]:=\int_{B_{0}}|\nabla v|_{g_{B}}^{2}\,d\mu_{B},\qquad\mathrm{Dom}(E_{B})=H^{1}_{0}(B_{0}).$
Total-space Dirichlet form.
Define
(19) $E_{t}[u]:=\int_{X_{t,0}}|\nabla u|_{g(t)}^{2}\,d\mu_{t},\qquad\mathrm{Dom}(E_{t})=H^{1}_{0}(X_{t,0}).$
Both forms are densely defined, closed, Markovian, and generate nonnegative selfadjoint operators $H_{B}$ and $H_{t}$ via the Friedrichs construction (we will use this in the next section); see [15, 16].
3.2. Normalized identifications and the fiber-constant projection
We recall the normalized disintegration and the lift/average operators from Section 2, but we record explicitly the normalization used in the present paper.
Assumption 3.1 (Normalized disintegration on $B_{0}$).
There exist measurable functions $A_{t}:B_{0}\to(0,\infty)$ and $\rho_{t}:X_{t,0}\to(0,\infty)$ such that for all $f\in L^{1}(X_{t,0},\mu_{t})$,
(20) $\int_{X_{t,0}}f\,d\mu_{t}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b),$
where $d\vartheta_{b,t}$ is the Haar probability measure on the flat torus $(F_{b,t},g_{F,t})$, and
(21) $\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1\qquad\text{for all }b\in B_{0}.$
Moreover, under the semi-flat $C^{1}$ control, we have the quantitative bounds
(22) $\|\rho_{t}-1\|_{L^{\infty}(X_{t,0})}+\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}+\|\nabla_{B}\rho_{t}\|_{L^{\infty}(X_{t,0})}\ \leq\ C\,\varepsilon_{t},$
with $\varepsilon_{t}\downarrow 0$. Here $\nabla_{B}\rho_{t}$ denotes the horizontal (base-direction) derivative with respect to the product splitting induced by the reference metric $g_{t}^{\Pi}=g_{B}\oplus g_{F,t}$ in local semi-flat coordinates.
Definition 3.2 (Normalized lift and fiber-average).
Define bounded linear maps $I_{t}:L^{2}_{B}\to L^{2}_{t}$ and $P_{t}:L^{2}_{t}\to L^{2}_{B}$ by
(23) $(I_{t}v)(x):=A_{t}(\Pi_{t}x)^{-1/2}\,v(\Pi_{t}x),\qquad(P_{t}u)(b):=A_{t}(b)^{1/2}\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}.$
Lemma 3.3 (Adjoint lock and projection).
The maps $(I_{t},P_{t})$ satisfy
(24) $P_{t}I_{t}=\mathrm{Id}_{L^{2}(B_{0})},\qquad I_{t}=P_{t}^{*}.$
In particular, the operator
(25) $\Pi_{t}^{\mathrm{fc}}:=I_{t}P_{t}:L^{2}_{t}\to L^{2}_{t}$
is the $L^{2}_{t}$-orthogonal projection onto the closed subspace of fiber-constant functions, and $(\Pi_{t}^{\mathrm{fc}})^{\perp}:=\mathrm{Id}-\Pi_{t}^{\mathrm{fc}}$ is the orthogonal projection onto its complement.
Proof.
The identity $P_{t}I_{t}=\mathrm{Id}$ follows immediately from (23) and (21). For adjointness, compute using (20): for $\phi\in L^{2}_{B}$ and $u\in L^{2}_{t}$,
$\langle I_{t}\phi,u\rangle_{L^{2}_{t}}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}A_{t}(b)^{-1/2}\phi(b)\,u\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{B}(b)=\int_{B_{0}}\phi(b)\,(P_{t}u)(b)\,d\mu_{B}(b)=\langle\phi,P_{t}u\rangle_{L^{2}_{B}}.$
Thus $I_{t}=P_{t}^{*}$, hence $\Pi_{t}^{\mathrm{fc}}=I_{t}P_{t}$ is an orthogonal projection. ∎
3.3. Vertical Poincaré inequality from the spectral gap
We now implement the first link in the chain:
[…]
$\Longrightarrow\text{suppression of nonconstant fiber modes at low energy}.$
Let $\nabla_{F}$ denote the vertical gradient with respect to the flat fiber metric $g_{F,t}$ (equivalently, the vertical part of the product metric $g_{t}^{\Pi}=g_{B}\oplus g_{F,t}$).
Lemma 3.4 (Fiberwise weighted Poincaré inequality).
Assume (22) and the spectral gap (6). Then there exists $C>0$ independent of $t$ and $b\in B_{0}$ such that for every $u\in H^{1}(F_{b,t})$,
(26) $\int_{F_{b,t}}\Big|u-\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}\Big|^{2}\rho_{t}\,d\vartheta_{b,t}\ \leq\ C\,s(t)^{2}\int_{F_{b,t}}|\nabla_{F}u|_{g_{F,t}}^{2}\,\rho_{t}\,d\vartheta_{b,t}.$
Proof.
Fix $b$. Let $m:=\int_{F_{b,t}}u\,d\vartheta_{b,t}$ and $m_{\rho}:=\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}$. Since $\int\rho_{t}\,d\vartheta_{b,t}=1$, we have
$m-m_{\rho}=\int_{F_{b,t}}(u-m)(1-\rho_{t})\,d\vartheta_{b,t},$
hence by Cauchy–Schwarz and $\|\rho_{t}-1\|_{\infty}\leq C\varepsilon_{t}$,
$|m-m_{\rho}|\leq\|\rho_{t}-1\|_{L^{2}(F_{b,t},d\vartheta)}\,\|u-m\|_{L^{2}(F_{b,t},d\vartheta)}\leq C\varepsilon_{t}\|u-m\|_{L^{2}(F_{b,t},d\vartheta)}.$
Therefore,
$\|u-m_{\rho}\|_{L^{2}(d\vartheta)}\leq\|u-m\|_{L^{2}(d\vartheta)}+|m-m_{\rho}|\leq(1+C\varepsilon_{t})\|u-m\|_{L^{2}(d\vartheta)}.$
By the standard Poincaré inequality on the flat torus and the gap (6),
$\|u-m\|_{L^{2}(d\vartheta)}^{2}\leq\lambda_{1}(F_{b,t},g_{F,t})^{-1}\,\|\nabla_{F}u\|_{L^{2}(d\vartheta)}^{2}\leq C\,s(t)^{2}\,\|\nabla_{F}u\|_{L^{2}(d\vartheta)}^{2}.$
Finally, since $\rho_{t}$ is uniformly comparable to 11 by (22), $\|\cdot\|_{L^{2}(\rho_{t}d\vartheta)}^{2}\simeq\|\cdot\|_{L^{2}(d\vartheta)}^{2}$ with constants $1+O(\varepsilon_{t})$, and similarly for the gradient term; absorbing these factors yields (26). ∎
Integrating fiberwise yields the global vertical Poincaré inequality for the orthogonal complement of the fiber-constant sector.
Proposition 3.5 (Global vertical Poincaré inequality).
Under the assumptions of Lemma 3.4, there exists $C>0$ independent of $t$ such that for all $u\in H^{1}(X_{t,0})$,
(27) $\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}\ \leq\ C\,s(t)^{2}\int_{X_{t,0}}|\nabla_{F}u|_{g_{F,t}}^{2}\,d\mu_{t}.$
Equivalently,
(28) $\int_{X_{t,0}}|\nabla_{F}u|_{g_{F,t}}^{2}\,d\mu_{t}\ \geq\ c\,s(t)^{-2}\,\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}.$
Proof.
By Lemma 3.3, $(\Pi_{t}^{\mathrm{fc}}u)|_{F_{b,t}}$ equals the fiberwise constant $m_{\rho}(b):=\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t}$ (pulled back to the fiber). Thus $((\Pi_{t}^{\mathrm{fc}})^{\perp}u)|_{F_{b,t}}=u-m_{\rho}(b)$. Apply Lemma 3.4 fiberwise and integrate over $b$ using the disintegration (20). This yields (27). The coercive form (28) is the same inequality rewritten. ∎
3.4. Energy comparison for lift and average
To reach Mosco convergence we need two quantitative compatibilities: (i) the lift $I_{t}$ is $H^{1}$-compatible (fiber-constants have the correct horizontal energy), (ii) the average $P_{t}$ does not increase the base energy beyond the total energy.
Lemma 3.6 (Energy of fiber-constant lifts).
There exists $C>0$ independent of $t$ such that for all $v\in H^{1}_{0}(B_{0})$,
(29) $E_{t}[I_{t}v]=E_{B}[v]\ +\ O(\varepsilon_{t})\,\|v\|_{H^{1}(B_{0})}^{2}.$
In particular, $\lim_{t\downarrow 0}E_{t}[I_{t}v]=E_{B}[v]$ for each fixed $v\in H^{1}_{0}(B_{0})$.
Proof.
This is the $H^{1}$-compatibility established in Section 2: since $I_{t}v=A_{t}^{-1/2}\,v\circ\Pi_{t}$, the vertical gradient vanishes identically and the horizontal gradient differs from $\nabla v$ only through $\nabla\log A_{t}$, which is $O(\varepsilon_{t})$ by (22). The $C^{1}$ semi-flat control ensures that the true metric $g(t)$ is uniformly close to the product metric on $X_{t,0}$, so replacing $g(t)$ by $g_{B}\oplus g_{F,t}$ introduces only an $O(\varepsilon_{t})$ relative error. Collecting terms yields (29). ∎
Lemma 3.7 (Energy contraction under averaging).
There exists $C>0$ independent of $t$ such that for all $u\in H^{1}_{0}(X_{t,0})$,
(30) $E_{B}[P_{t}u]\ \leq\ (1+C\varepsilon_{t})\,E_{t}[u]\ +\ C\varepsilon_{t}\,\|u\|_{L^{2}_{t}}^{2}.$
In particular, if $\sup_{t}\|u_{t}\|_{L^{2}_{t}}<\infty$ and $\sup_{t}E_{t}[u_{t}]<\infty$, then $\{P_{t}u_{t}\}$ is bounded in $H^{1}_{0}(B_{0})$.
Proof.
We sketch the argument, which is standard once one has the quantitative disintegration and the $C^{1}$ semi-flat control.
Write $v:=P_{t}u$. In local semi-flat coordinates on $X_{t,0}\simeq B_{0}\times\mathbb{T}^{2}$,
$v(b)=A_{t}(b)^{1/2}\int_{\mathbb{T}^{2}}u(b,\theta)\,\rho_{t}(b,\theta)\,d\vartheta.$
Differentiate under the integral sign along a base vector field and use the product splitting to interpret $\nabla_{B}u$ as the horizontal derivative. The leading term is $A_{t}^{1/2}\int\nabla_{B}u\ \rho_{t}\,d\vartheta$, which is controlled by Jensen:
$\Big|\int\nabla_{B}u\ \rho_{t}\,d\vartheta\Big|^{2}\leq\int|\nabla_{B}u|^{2}\,\rho_{t}\,d\vartheta.$
All remaining terms come from differentiating $A_{t}^{1/2}$ and $\rho_{t}$ and are therefore bounded by $\|\nabla\log A_{t}\|_{\infty}$ and $\|\nabla_{B}\rho_{t}\|_{\infty}$, hence by $C\varepsilon_{t}$ via (22). These error terms are controlled by Cauchy–Schwarz, yielding a contribution bounded by $C\varepsilon_{t}\big(\|\nabla_{B}u\|_{L^{2}_{t}}^{2}+\|u\|_{L^{2}_{t}}^{2}\big)$. Finally, since $|\nabla_{B}u|^{2}\leq|\nabla u|_{g(t)}^{2}$ up to $O(\varepsilon_{t})$ errors by the $C^{1}$ semi-flat control, we obtain (30). The Dirichlet boundary condition implies $v|_{\partial B_{0}}=0$, hence $v\in H^{1}_{0}(B_{0})$. ∎
3.5. Mosco convergence of Dirichlet forms on identified spaces
We now prove the second link in the chain:
$\text{form comparison + $H^{1}$-compatibility of }(I_{t},P_{t})\Longrightarrow\text{Mosco convergence}.$
Definition 3.8 (Mosco convergence via $(P_{t},I_{t})$).
We say that $E_{t}$ Mosco-converges to $E_{B}$ along $(P_{t},I_{t})$ in the sense of Mosco; see [13, 14, 6] if:
  • (M1) Liminf:
    whenever $u_{t}\in H^{1}_{0}(X_{t,0})$ satisfy $\sup_{t}\|u_{t}\|_{L^{2}_{t}}<\infty$ and $P_{t}u_{t}\to v$ in $L^{2}_{B}$, one has $v\in H^{1}_{0}(B_{0})$ and
    $\liminf_{t\downarrow 0}E_{t}[u_{t}]\ \geq\ E_{B}[v];$
  • (M2) Recovery:
    for every $v\in H^{1}_{0}(B_{0})$ there exist $u_{t}\in H^{1}_{0}(X_{t,0})$ such that $P_{t}u_{t}=v$ and
    $\limsup_{t\downarrow 0}E_{t}[u_{t}]\ \leq\ E_{B}[v].$
Theorem 3.9 (Mosco convergence).
Assume the semi-flat $C^{1}$ control on $X_{t,0}$ (hence (22)) and the uniform vertical gap Assumption $2.$ Then the Dirichlet forms $E_{t}$ Mosco-converge to $E_{B}$ on $B_{0}$ in the sense of Definition 3.8 via the identification pair $(P_{t},I_{t})$.
Proof.
Step 1: (M1) Liminf. Let $u_{t}\in H^{1}_{0}(X_{t,0})$ with $\sup_{t}\|u_{t}\|_{L^{2}_{t}}<\infty$ and $P_{t}u_{t}\to v$ in $L^{2}_{B}$. If $\liminf_{t}E_{t}[u_{t}]=\infty$ there is nothing to prove, so assume $\sup_{t}E_{t}[u_{t}]<\infty$ along a subsequence.
By Lemma 3.7, the sequence $v_{t}:=P_{t}u_{t}$ is bounded in $H^{1}_{0}(B_{0})$, hence (up to a subsequence) $v_{t}\rightharpoonup v$ weakly in $H^{1}_{0}(B_{0})$ and $v_{t}\to v$ strongly in $L^{2}(B_{0})$. By weak lower semicontinuity,
(31) $E_{B}[v]\ \leq\ \liminf_{t\downarrow 0}E_{B}[v_{t}].$
On the other hand, (30) gives
$E_{B}[v_{t}]\ \leq\ (1+C\varepsilon_{t})E_{t}[u_{t}]+C\varepsilon_{t}\|u_{t}\|_{L^{2}_{t}}^{2}.$
Taking $\liminf$ and using $\varepsilon_{t}\to 0$ and $\sup_{t}\|u_{t}\|_{L^{2}_{t}}<\infty$ yields
(32) $\liminf_{t\downarrow 0}E_{B}[v_{t}]\ \leq\ \liminf_{t\downarrow 0}E_{t}[u_{t}].$
Combining (31) and (32) proves (M1).
Step 2: (M2) Recovery. Fix $v\in H^{1}_{0}(B_{0})$ and set $u_{t}:=I_{t}v$. Then $u_{t}\in H^{1}_{0}(X_{t,0})$ and $P_{t}u_{t}=P_{t}I_{t}v=v$ by Lemma $3.$ By Lemma 3.6,
$\limsup_{t\downarrow 0}E_{t}[u_{t}]=\limsup_{t\downarrow 0}E_{t}[I_{t}v]\leq E_{B}[v],$
which proves (M2). ∎
The vertical gap also yields an explicit suppression of the fiber-oscillatory component along bounded-energy sequences, which will be used in the semigroup/kernels step.
Corollary 3.10 (Vanishing of fiber-oscillatory mass at bounded energy).
Assume the hypotheses of Theorem 3.9. If $u_{t}\in H^{1}_{0}(X_{t,0})$ satisfy $\sup_{t}E_{t}[u_{t}]<\infty$, then
$\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u_{t}\|_{L^{2}_{t}}\ \longrightarrow\ 0\qquad(t\downarrow 0).$
Proof.
By Proposition 3.5 and $|\nabla_{F}u|^{2}\leq|\nabla u|_{g(t)}^{2}$ up to $O(\varepsilon_{t})$,
$\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u_{t}\|_{L^{2}_{t}}^{2}\leq Cs(t)^{2}\int_{X_{t,0}}|\nabla_{F}u_{t}|^{2}\,d\mu_{t}\leq Cs(t)^{2}\,E_{t}[u_{t}].$
Since $s(t)\downarrow 0$ and $\sup_{t}E_{t}[u_{t}]<\infty$, the right-hand side tends to 0. ∎
3.6. Section summary
We have proved: (i) the vertical gap yields a global vertical Poincaré inequality for $(\Pi_{t}^{\mathrm{fc}})^{\perp}$, hence suppression of nonconstant fiber modes at bounded energy; (ii) the lift and average maps $(I_{t},P_{t})$ are $H^{1}$-compatible in the precise form needed to establish Mosco convergence of the Dirichlet forms $E_{t}\to E_{B}$ on $B_{0}$.
In the next section we apply the standard implication
$\text{Mosco convergence}\ \Longrightarrow\ \text{strong resolvent/semigroup convergence}$
for the associated selfadjoint operators, and then convert semigroup convergence into the interior bilinear heat-kernel pairing limit via kernel representation.
4. From Mosco convergence to compressed semigroups and the interior bilinear kernel limit
In this section we close the main interior chain:
$\text{Mosco convergence of }E_{t}$ $\Longrightarrow\text{strong resolvent/semigroup convergence}$
$\Longrightarrow\text{bilinear heat-kernel pairing limit}$
$\qquad\text{on }B_{0}\Subset B_{\mathrm{reg}}.$
Throughout, $B_{0}\Subset B_{\mathrm{reg}}$ is fixed and $X_{t,0}:=\Pi_{t}^{-1}(B_{0})$. We work under Dirichlet boundary conditions; the Neumann and nonnegatively weighted Robin cases are treated later and do not affect the arguments of the present section.
4.1. Generators and the compressed operators
Let $H_{t}$ and $H_{B}$ denote the nonnegative selfadjoint operators associated with the closed forms $E_{t}$ and $E_{B}$ from (19)–(18), i.e. the Friedrichs realizations of $\Delta_{g(t)}$ on $X_{t,0}$ and $\Delta_{g_{B}}$ on $B_{0}$, respectively.
Write the heat semigroups
$T_{t}(\tau):=e^{-\tau H_{t}}\quad\text{on }L^{2}_{t},\qquad T_{B}(\tau):=e^{-\tau H_{B}}\quad\text{on }L^{2}_{B},\qquad\tau>0.$
Both are selfadjoint, positivity preserving, and $L^{2}$-contractive.
We define the compressed (fiber-constant) heat operators on the base Hilbert space $L^{2}_{B}$ by
(33) $S_{t}(\tau):=P_{t}T_{t}(\tau)I_{t}\;=\;I_{t}^{*}\,T_{t}(\tau)\,I_{t}:L^{2}_{B}\to L^{2}_{B}.$
By Lemma 3.3 and contractivity of $T_{t}(\tau)$, each $S_{t}(\tau)$ is a selfadjoint contraction on $L^{2}_{B}$:
(34) $\|S_{t}(\tau)\|_{L^{2}_{B}\to L^{2}_{B}}\leq 1,\qquad S_{t}(\tau)=S_{t}(\tau)^{*}.$
4.2. Mosco convergence implies strong resolvent convergence
We first record the standard implication “Mosco $\Rightarrow$ strong resolvent convergence” in our identified setting. For completeness we include a short variational proof, adapted to the connecting maps $(P_{t},I_{t})$.
Theorem 4.1 (Strong resolvent convergence via Mosco).
Assume $E_{t}\overset{M}{\longrightarrow}E_{B}$ via $(P_{t},I_{t})$ in the sense of Theorem 3.9. Then for every $\alpha>0$ and every $f\in L^{2}_{B}$,
(35) $P_{t}(\alpha+H_{t})^{-1}I_{t}f\longrightarrow(\alpha+H_{B})^{-1}f\qquad\text{strongly in }L^{2}_{B}.$
Equivalently,
$P_{t}(\alpha+H_{t})^{-1}I_{t}\;\xrightarrow[t\downarrow 0]{\ \mathrm{s}\ }\;(\alpha+H_{B})^{-1}\quad\text{in }\mathcal{B}(L^{2}_{B}).$
Proof.
Fix $\alpha>0$ and $f\in L^{2}_{B}$. Consider the strictly convex functionals
$\mathcal{J}_{t}[u]:=E_{t}[u]+\alpha\|u\|_{L^{2}_{t}}^{2}-2\alpha\langle u,I_{t}f\rangle_{L^{2}_{t}},\qquad u\in H^{1}_{0}(X_{t,0}),$
and
$\mathcal{J}_{B}[v]:=E_{B}[v]+\alpha\|v\|_{L^{2}_{B}}^{2}-2\alpha\langle v,f\rangle_{L^{2}_{B}},\qquad v\in H^{1}_{0}(B_{0}).$
By standard form theory, $\mathcal{J}_{t}$ and $\mathcal{J}_{B}$ have unique minimizers
$u_{t}=(\alpha+H_{t})^{-1}I_{t}f\in H^{1}_{0}(X_{t,0}),\qquad v=(\alpha+H_{B})^{-1}f\in H^{1}_{0}(B_{0}).$
Step 1: compactness of the pushed-forward minimizers. Set $v_{t}:=P_{t}u_{t}\in L^{2}_{B}$. Since $u_{t}$ minimizes $\mathcal{J}_{t}$, we have
$\mathcal{J}_{t}[u_{t}]\leq\mathcal{J}_{t}[I_{t}v]=E_{t}[I_{t}v]+\alpha\|I_{t}v\|_{L^{2}_{t}}^{2}-2\alpha\langle I_{t}v,I_{t}f\rangle_{L^{2}_{t}}.$
Using $I_{t}=P_{t}^{*}$ and $P_{t}I_{t}=\mathrm{Id}$, we have the exact identities
$\|I_{t}v\|_{L^{2}_{t}}^{2}=\langle I_{t}v,I_{t}v\rangle_{L^{2}_{t}}=\langle v,P_{t}I_{t}v\rangle_{L^{2}_{B}}=\|v\|_{L^{2}_{B}}^{2},$
and
$\langle I_{t}v,I_{t}f\rangle_{L^{2}_{t}}=\langle v,P_{t}I_{t}f\rangle_{L^{2}_{B}}=\langle v,f\rangle_{L^{2}_{B}}.$
Moreover, Lemma 3.6 gives $E_{t}[I_{t}v]=E_{B}[v]+o(1)$. Hence $\sup_{t}\mathcal{J}_{t}[u_{t}]<\infty$, which implies $\sup_{t}\|u_{t}\|_{L^{2}_{t}}<\infty$ and $\sup_{t}E_{t}[u_{t}]<\infty$. By Lemma 3.7, $\{v_{t}\}$ is bounded in $H^{1}_{0}(B_{0})$, thus (up to a subsequence) $v_{t}\rightharpoonup v_{\infty}$ weakly in $H^{1}_{0}$ and $v_{t}\to v_{\infty}$ strongly in $L^{2}_{B}$.
Step 2: identify the limit as the base resolvent. By Mosco (M1), applied to $u_{t}$ with $P_{t}u_{t}=v_{t}\to v_{\infty}$ in $L^{2}_{B}$,
$\liminf_{t\downarrow 0}E_{t}[u_{t}]\geq E_{B}[v_{\infty}].$
Also, using $I_{t}=P_{t}^{*}$ and $P_{t}u_{t}=v_{t}$,
$\langle u_{t},I_{t}f\rangle_{L^{2}_{t}}=\langle P_{t}u_{t},f\rangle_{L^{2}_{B}}=\langle v_{t},f\rangle_{L^{2}_{B}}\to\langle v_{\infty},f\rangle_{L^{2}_{B}}.$
Finally, since $\|P_{t}\|=\|I_{t}\|=1$ (as $I_{t}$ is an isometry and $P_{t}=I_{t}^{*}$), we have
$\|u_{t}\|_{L^{2}_{t}}\geq\|P_{t}u_{t}\|_{L^{2}_{B}}=\|v_{t}\|_{L^{2}_{B}},$
hence
$\liminf_{t\downarrow 0}\|u_{t}\|_{L^{2}_{t}}^{2}\geq\lim_{t\downarrow 0}\|v_{t}\|_{L^{2}_{B}}^{2}=\|v_{\infty}\|_{L^{2}_{B}}^{2}.$
Putting these together yields
$\liminf_{t\downarrow 0}\mathcal{J}_{t}[u_{t}]\ \geq\ \mathcal{J}_{B}[v_{\infty}].$
On the other hand, the recovery sequence $I_{t}v$ (Mosco (M2)) implies
$\limsup_{t\downarrow 0}\inf\mathcal{J}_{t}\ \leq\ \limsup_{t\downarrow 0}\mathcal{J}_{t}[I_{t}v]\ =\ \mathcal{J}_{B}[v].$
Therefore $\mathcal{J}_{B}[v_{\infty}]\leq\mathcal{J}_{B}[v]$. By uniqueness of the minimizer of $\mathcal{J}_{B}$, we have $v_{\infty}=v$. Since the limit is unique, the whole family $v_{t}=P_{t}u_{t}$ converges to $v$ in $L^{2}_{B}$, proving (35). ∎
4.3. Strong convergence of compressed heat operators
We now pass from resolvents to heat operators.
Theorem 4.2 (Strong convergence of compressed heat operators).
Assume the hypotheses of Theorem 4.1. Then for every $\tau>0$,
(36) $S_{t}(\tau)=P_{t}e^{-\tau H_{t}}I_{t}\longrightarrow e^{-\tau H_{B}}=T_{B}(\tau)\qquad\text{strongly in }\mathcal{B}(L^{2}_{B}).$
Moreover, the convergence is uniform for $\tau$ in compact subsets of $(0,\infty)$.
Proof.
For nonnegative selfadjoint operators, strong resolvent convergence implies strong convergence of the bounded functional calculus for every bounded continuous function vanishing at infinity [17, 18]. In particular, for each fixed $\tau>0$ the function $x\mapsto e^{-\tau x}$ yields
$P_{t}e^{-\tau H_{t}}I_{t}\to e^{-\tau H_{B}}\quad\text{strongly on }L^{2}_{B}.$
A direct proof can be obtained by approximating $e^{-\tau x}$ uniformly on $0,\infty)$ by rational functions in $(\alpha+x)^{-1}$ (or by Euler approximants $(1+\tau x/n)^{-n}$) and using Theorem 4.1 together with uniform operator bounds (34). Uniformity on compact $\tau$-intervals follows from uniform approximation. ∎
4.4. Heat kernels and bilinearization
On the precompact domains $B_{0}$ and $X_{t,0}$ (with smooth boundary), the resolvents are compact and the heat operators admit integral kernels.
Lemma 4.3 (Kernel representation).
For each $\tau>0$ there exist symmetric kernels
$K_{t}(\cdot,\cdot;\tau)\in L^{2}(X_{t,0}\times X_{t,0}),\qquad K_{B}(\cdot,\cdot;\tau)\in L^{2}(B_{0}\times B_{0}),$
such that for all $F,G\in L^{2}_{t}$ and all $\phi,\psi\in L^{2}_{B}$,
(37) $\langle F,T_{t}(\tau)G\rangle_{L^{2}_{t}}$ $=\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,F(x)\,G(y)\,d\mu_{t}(x)\,d\mu_{t}(y),$
(38) $\langle\phi,T_{B}(\tau)\psi\rangle_{L^{2}_{B}}$ $=\int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\phi(b)\,\psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
Proof sketch.
On a precompact domain with Dirichlet boundary, the embedding $H^{1}_{0}\hookrightarrow L^{2}$ is compact [20], hence $(\alpha+H)^{-1}$ is compact and $H$ has discrete spectrum with an $L^{2}$-orthonormal basis of eigenfunctions. The kernel is given by the usual spectral expansion $K(x,y;\tau)=\sum_{k}e^{-\tau\lambda_{k}}\varphi_{k}(x)\varphi_{k}(y)$, converging in $L^{2}$ on the product domain; see [7, 8]
The compressed operator $S_{t}(\tau)$ admits an exact bilinear kernel expression.
Proposition 4.4 (Bilinearization identity).
For every $\tau>0$ and every $\Phi,\Psi\in L^{2}_{B}$,
(39) $\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\langle I_{t}\Phi,T_{t}(\tau)I_{t}\Psi\rangle_{L^{2}_{t}}.$
If the kernels of Lemma 4.3 are used, then
(40) $\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y),$
and similarly
(41) $\langle\Phi,T_{B}(\tau)\Psi\rangle_{L^{2}_{B}}=\int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
Proof.
By Lemma 3.3, $I_{t}=P_{t}^{*}$, hence
$\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}=\langle\Phi,P_{t}T_{t}(\tau)I_{t}\Psi\rangle_{L^{2}_{B}}=\langle I_{t}\Phi,T_{t}(\tau)I_{t}\Psi\rangle_{L^{2}_{t}},$
which is (39). The kernel identities follow from (37) and (38). ∎
4.5. Interior bilinear heat-kernel pairing limit
We can now state the interior bilinear limit as a direct corollary of Theorem 4.2.
Theorem 4.5 (Interior bilinear heat-kernel pairing limit).
Assume the hypotheses of Theorem 4.2, and let $B_{0}\Subset B_{\mathrm{reg}}$. Then for every $\tau>0$ and every $\Phi,\Psi\in L^{2}(B_{0})$,
(42) $\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y)\ \longrightarrow\ \int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
In particular, the convergence holds for all $\Phi,\Psi\in C_{c}^{\infty}(B_{0})$.
Proof.
By Proposition 4.4, the left-hand side of (42) equals $\langle\Phi,S_{t}(\tau)\Psi\rangle_{L^{2}_{B}}$ and the right-hand side equals $\langle\Phi,T_{B}(\tau)\Psi\rangle_{L^{2}_{B}}$. By Theorem 4.2, $S_{t}(\tau)\Psi\to T_{B}(\tau)\Psi$ strongly in $L^{2}_{B}$; testing against $\Phi$ gives (42). ∎
4.6. Strong convergence versus bilinear pairing limits (a rigorous equivalence)
For later use (and to match the “compressed semigroup $\Leftrightarrow$ bilinear kernel” viewpoint precisely), we record a rigorous equivalence at the level of the whole family $\{\tau>0\}$. The only extra input needed beyond bilinear matrix-coefficient convergence is an “asymptotic semigroup” property for $S_{t}(\tau)$, which follows from the vertical gap and smoothing.
Lemma 4.6 (Energy smoothing for the heat flow).
For every $\sigma>0$ and every $G\in L^{2}_{t}$,
(43) $E_{t}[T_{t}(\sigma)G]\ \leq\ \frac{C}{\sigma}\,\|G\|_{L^{2}_{t}}^{2},$
where $C>0$ is universal.
Proof.
By spectral calculus,
$E_{t}[T_{t}(\sigma)G]=\|H_{t}^{1/2}e^{-\sigma H_{t}}G\|_{L^{2}_{t}}^{2}=\int_{0}^{\infty}\lambda\,e^{-2\sigma\lambda}\,d\|E_{\lambda}G\|^{2}\leq\Big(\sup_{\lambda\geq 0}\lambda e^{-2\sigma\lambda}\Big)\|G\|_{L^{2}_{t}}^{2}\leq\frac{C}{\sigma}\|G\|_{L^{2}_{t}}^{2}.$
Lemma 4.7 (Leakage estimate from fiber-constant data).
Assume the vertical Poincaré inequality (27) from Proposition 3.5. Then for every $\sigma>0$ and every $f\in L^{2}_{B}$,
(44) $\|(\Pi_{t}^{\mathrm{fc}})^{\perp}\,T_{t}(\sigma)I_{t}f\|_{L^{2}_{t}}\ \leq\ C\,\frac{s(t)}{\sqrt{\sigma}}\;\|f\|_{L^{2}_{B}},$
with $C$ independent of $t$.
Proof.
Apply (27) to $u=T_{t}(\sigma)I_{t}f$ and use $|\nabla_{F}u|^{2}\leq|\nabla u|^{2}$:
$\|(\Pi_{t}^{\mathrm{fc}})^{\perp}u\|_{L^{2}_{t}}^{2}\leq Cs(t)^{2}\int_{X_{t,0}}|\nabla u|^{2}\,d\mu_{t}=Cs(t)^{2}\,E_{t}[u].$
Then invoke Lemma 4.6 and the $L^{2}$-isometry of $I_{t}$ to get $E_{t}[u]\leq C\sigma^{-1}\|I_{t}f\|_{L^{2}_{t}}^{2}=C\sigma^{-1}\|f\|_{L^{2}_{B}}^{2}$. ∎
Proposition 4.8 (Asymptotic semigroup property of $S_{t}(\tau)$).
For all $\tau,\sigma>0$,
(45) $\|S_{t}(\tau)S_{t}(\sigma)-S_{t}(\tau+\sigma)\|_{L^{2}_{B}\to L^{2}_{B}}\ \leq\ C\,\frac{s(t)}{\sqrt{\sigma}}.$
In particular, for each fixed $\tau>0$, $\|S_{t}(\tau)^{2}-S_{t}(2\tau)\|_{op}\to 0$ as $t\downarrow 0$.
Proof.
Using $P_{t}I_{t}=\mathrm{Id}$ and $\Pi_{t}^{\mathrm{fc}}=I_{t}P_{t}$,
$S_{t}(\tau)S_{t}(\sigma)-S_{t}(\tau+\sigma)=P_{t}T_{t}(\tau)I_{t}P_{t}T_{t}(\sigma)I_{t}-P_{t}T_{t}(\tau)T_{t}(\sigma)I_{t}=-P_{t}T_{t}(\tau)(\Pi_{t}^{\mathrm{fc}})^{\perp}T_{t}(\sigma)I_{t}.$
Taking operator norms and using $\|P_{t}\|\leq 1$, $\|T_{t}(\tau)\|\leq 1$ and Lemma 4.7 gives (45). ∎
We can now state a rigorous “strong $\Leftrightarrow$ bilinear kernel” equivalence for the full time family.
Theorem 4.9 (Strong compressed limit $\Longleftrightarrow$ bilinear kernel limits).
Assume Proposition 4.8. Then the following two statements are equivalent:
  • (S)
    For every $\tau>0$, $S_{t}(\tau)\to T_{B}(\tau)$ strongly on $L^{2}_{B}$.
  • (K)
    For every $\tau>0$ and every $\Phi,\Psi\in L^{2}(B_{0})$,
    (46) $\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y)\ \longrightarrow\ \int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
Proof.
[…]
To upgrade to strong convergence, it suffices to show $\|x_{t}\|\to\|T_{B}(\tau)f\|$. Since $S_{t}(\tau)$ is selfadjoint,
$\|x_{t}\|_{L^{2}_{B}}^{2}=\langle S_{t}(\tau)f,S_{t}(\tau)f\rangle=\langle f,S_{t}(\tau)^{2}f\rangle.$
By Proposition 4.8,
$\langle f,S_{t}(\tau)^{2}f\rangle=\langle f,S_{t}(2\tau)f\rangle+o(1)\|f\|^{2}.$
Applying (K) at time $2\tau$ yields $\langle f,S_{t}(2\tau)f\rangle\to\langle f,T_{B}(2\tau)f\rangle$. Finally, since $T_{B}(\tau)$ is a selfadjoint semigroup,
$\langle f,T_{B}(2\tau)f\rangle=\langle T_{B}(\tau)f,T_{B}(\tau)f\rangle=\|T_{B}(\tau)f\|^{2}.$
Thus $\|x_{t}\|^{2}\to\|T_{B}(\tau)f\|^{2}$, and together with weak convergence this implies $x_{t}\to T_{B}(\tau)f$ strongly. Since $f$ was arbitrary, we obtain (S). ∎
4.7. Section summary
We have shown the following: (i) Mosco convergence of $E_{t}$ to $E_{B}$ implies strong resolvent convergence (35), hence strong convergence of the compressed heat operators $S_{t}(\tau)\to T_{B}(\tau)$ for every $\tau>0$; (ii) by kernel representation and bilinearization, strong convergence yields the interior bilinear heat-kernel pairing limit (42) on $B_{0}\Subset B_{\mathrm{reg}}$; and (iii) conversely, the family of bilinear kernel limits for all $\tau>0$ implies strong convergence of $S_{t}(\tau)$, once the asymptotic semigroup property (45) is observed, a property that follows from the vertical gap and smoothing.
In the next section we move beyond $B_{\mathrm{reg}}$: near each discriminant point $p_{j}\in D$ the base metric is modeled by a flat cone, and the bilinear pairing requires an intrinsic conic renormalization.
5. Conic parametrix on the base and the renormalized bilinear functional
Let $B$ be a smooth connected surface, and let
$D=\{p_{1},\dots,p_{N}\}\subset B$
be a finite discriminant set. Set $B_{\mathrm{reg}}:=B\setminus D$ and let $g_{B}$ be a smooth Riemannian metric on $B_{\mathrm{reg}}$. We assume that $g_{B}$ has isolated conic asymptotics at each $p_{j}$ in the sense below.
This section has four goals:
  1. (1)
    Build wedge charts near $D$ and formulate a $C^{1}$ conic approximation of $g_{B}$.
  2. (2)
    Introduce the flat-cone heat kernel and construct a conic parametrix for the base heat kernel with weighted remainders.
  3. (3)
    Define a cutoff-independent renormalized bilinear functional $K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)$.
  4. (4)
    Identify the collapsed total-space bilinear limits with $K^{\mathrm{ren}}_{B}$ for each fixed $\tau>0$.
5.1. Wedge charts and $C^{1}$ conic asymptotics of $g_{B}$ near $D$
Fix $p_{j}\in D$. Let $U_{j}\subset B$ be a coordinate neighborhood of $p_{j}$, and choose polar coordinates
$(r,\theta)\in(0,r_{0})\times\mathbb{S}^{1}\qquad\text{on }U_{j}\setminus\{p_{j}\}$
centered at $p_{j}$ (so that $r\downarrow 0$ corresponds to approaching $p_{j}$). We assume:
Assumption 5.1 ($C^{1}$ conic asymptotics).
There exist parameters $\alpha_{j}>0$ (cone angles $2\pi\alpha_{j}$) and $\beta\in(0,1]$ such that on $U_{j}\setminus\{p_{j}\}$,
(47) $g_{B}=g_{j}^{\mathrm{cone}}+q_{j},\qquad g_{j}^{\mathrm{cone}}:=dr^{2}+\alpha_{j}^{2}r^{2}\,d\theta^{2},$
where the $2$-tensor $q_{j}$ satisfies the weighted $C^{1}$ bounds
(48) $|q_{j}|_{g^{\mathrm{cone}}_{j}}\leq Cr^{\beta},\qquad|\nabla^{\mathrm{cone}}q_{j}|_{g^{\mathrm{cone}}_{j}}\leq Cr^{\beta-1},$
for some $C>0$ independent of $(r,\theta)$. In particular, the Riemannian measures satisfy
(49) $d\mu_{g_{B}}=(1+O(r^{\beta}))\,d\mu_{g^{\mathrm{cone}}_{j}},\qquad d\mu_{g^{\mathrm{cone}}_{j}}=\alpha_{j}r\,dr\,d\theta,$
and the distances induced by $g_{B}$ and $g^{\mathrm{cone}}_{j}$ are locally bi-Lipschitz equivalent on $U_{j}\setminus\{p_{j}\}$.
Remark 5.2.
Assumption 5.1 is precisely what we need for a stable conic heat parametrix with remainders controlled by $(r+r^{\prime})^{\beta}$ weights. It does not require $g_{B}$ to extend smoothly across $p_{j}$.
See Figure 2 for the wedge chart and the inner/outer decomposition used in the renormalization.
[…]
Figure 2. Local wedge chart near a discriminant point $p_{j}\in D$ and the corresponding flat-cone model. Panel (a) shows polar coordinates $(r,\theta)$ on $U_{j}\setminus\{p_{j}\}\subset B_{\mathrm{reg}}$, together with the inner region $r<\rho$ and outer region $r>\rho$ used in the renormalization. Panel (b) depicts the model cone $\mathsf{C}_{\alpha_{j}}$ with cone angle $2\pi\alpha_{j}$, for which $g_{B}$ is $C^{1}$-close in Assumption 5.1.
5.2. The flat-cone heat kernel and basic bounds
Let $\mathsf{C}_{\alpha_{j}}:=((0,\infty)\times\mathbb{S}^{1},\,g^{\mathrm{cone}}_{j})$ denote the flat cone of angle $2\pi\alpha_{j}$. Let $H^{\mathrm{cone}}_{j}$ be the Friedrichs realization of $\Delta_{g^{\mathrm{cone}}_{j}}$ on $L^{2}(\mathsf{C}_{\alpha_{j}},d\mu_{g^{\mathrm{cone}}_{j}})$. Denote by $K^{\mathrm{cone}}_{j}(\cdot,\cdot;\tau)$ its heat kernel.
One may write $K^{\mathrm{cone}}_{j}$ explicitly via Bessel functions; for our purposes we only need: positivity, symmetry, Gaussian domination, and scale-invariant derivative bounds; see [5, 9].
Lemma 5.3 (Model cone kernel estimates).
For each $j$ and each $\tau_{0}>0$ there exist constants $C,c>0$ (depending only on $\alpha_{j}$ and $\tau_{0}$) such that for all $0<\tau\leq\tau_{0}$ and all $(r,\theta),(r^{\prime},\theta^{\prime})\in(0,r_{0})\times\mathbb{S}^{1}$,
(50) $0\leq K^{\mathrm{cone}}_{j}((r,\theta),(r^{\prime},\theta^{\prime});\tau)\leq\frac{C}{\tau}\exp\!\Big(-\frac{d_{\mathrm{cone}}((r,\theta),(r^{\prime},\theta^{\prime}))^{2}}{c\,\tau}\Big),$
and
(51) $|\partial_{r}K^{\mathrm{cone}}_{j}|+|\partial_{r^{\prime}}K^{\mathrm{cone}}_{j}|\leq\frac{C}{\tau^{3/2}}\exp\!\Big(-\frac{d_{\mathrm{cone}}((r,\theta),(r^{\prime},\theta^{\prime}))^{2}}{c\,\tau}\Big).$
Remark 5.4.
Such bounds are standard for heat kernels on flat cones; they follow from the explicit Bessel expansion and from general heat kernel estimates on conic manifolds (e.g. the parametrix calculus).
5.3. Conic parametrix for the base heat kernel with weighted remainders
Let $H_{B}$ denote the Friedrichs realization of $\Delta_{g_{B}}$ on $L^{2}(B_{\mathrm{reg}},d\mu_{g_{B}})$, and let $K_{B}(\cdot,\cdot;\tau)$ be its heat kernel on $B_{\mathrm{reg}}\times B_{\mathrm{reg}}$.
Fix cutoffs $\chi_{j},\widetilde{\chi}_{j}\in C_{c}^{\infty}(U_{j})$ such that
$\chi_{j}\equiv 1\ \text{on a neighborhood of }p_{j},\qquad\widetilde{\chi}_{j}\equiv 1\ \text{on }\mathrm{supp}\,\chi_{j}.$
We view $K^{\mathrm{cone}}_{j}$ as a kernel on $U_{j}\setminus\{p_{j}\}$ via the wedge coordinates.
Proposition 5.5 (Local conic parametrix with $(r+r^{\prime})^{\beta}$ remainder).
Assume Assumption 5.1. A local conic heat parametrix with weighted remainders can be constructed by standard edge methods; see [9, 22, 5]. Then there exist $\tau_{0}\in(0,1]$ and kernels $R_{j}(\cdot,\cdot;\tau)$ on $(U_{j}\setminus\{p_{j}\})\times(U_{j}\setminus\{p_{j}\})$ such that for all $0<\tau\leq\tau_{0}$ and all $b,b^{\prime}\in U_{j}\setminus\{p_{j}\}$,
(52) $K_{B}(b,b^{\prime};\tau)=\chi_{j}(b)\chi_{j}(b^{\prime})\,K^{\mathrm{cone}}_{j}(b,b^{\prime};\tau)+R_{j}(b,b^{\prime};\tau),$
and the remainder satisfies the weighted bound
(53) $|R_{j}(b,b^{\prime};\tau)|\leq\frac{C}{\tau}\,(r(b)+r(b^{\prime}))^{\beta}\,\exp\!\Big(-\frac{d_{g_{B}}(b,b^{\prime})^{2}}{C\,\tau}\Big),\qquad 0<\tau\leq\tau_{0},$
where $r(\cdot)$ is the wedge radial variable around $p_{j}$.
Proof sketch.
Let $K^{(0)}:=\widetilde{\chi}_{j}\otimes\widetilde{\chi}_{j}\,K^{\mathrm{cone}}_{j}$. Using (47)–(48), the operator difference $H_{B}-H^{\mathrm{cone}}_{j}$ has coefficients of size $O(r^{\beta})$ and first derivatives of size $O(r^{\beta-1})$ in the wedge coordinates. A Duhamel parametrix construction yields
$(\partial_{\tau}+H_{B})K^{(0)}=\delta+\mathcal{E},$
where $\mathcal{E}$ is supported where $\nabla\widetilde{\chi}_{j}\neq 0$ or where $q_{j}\neq 0$, and satisfies Gaussian domination with a prefactor $(r+r^{\prime})^{\beta}\tau^{-1}$ thanks to (50)–(51) and (48). Setting
$R_{j}:=\int_{0}^{\tau}e^{-(\tau-s)H_{B}}\,\mathcal{E}(s)\,ds$
and using Gaussian bounds for $e^{-tH_{B}}$ on $B_{\mathrm{reg}}$ gives (53). ∎
5.4. Definition of the renormalized bilinear functional
Let $\tau>0$ and let $\Phi,\Psi\in C_{c}^{\infty}(B)$ be smooth test functions, not assumed to avoid $D$. (We identify $\Phi,\Psi$ with their restrictions to $B_{\mathrm{reg}}$ whenever needed.)
Radial cutoff at scale $\rho$.
Fix a smooth function $\eta\in C^{\infty}([0,\infty))$ with $\eta\equiv 0$ on $[0,1/2]$ and $\eta\equiv 1$ on $1,\infty)$. For $\rho\in(0,r_{0})$ define $\eta_{\rho}(r):=\eta(r/\rho)$. Using wedge charts near each $p_{j}$, define a global cutoff $\eta_{\rho}\in C^{\infty}(B)$ by
$\eta_{\rho}(b):=1-\sum_{j=1}^{N}(1-\eta_{\rho}(r_{j}(b)))\,\chi_{j}(b),$
where $r_{j}(\cdot)$ is the wedge radial coordinate near $p_{j}$ and $\chi_{j}$ are the cutoffs in the parametrix. Since $D$ is finite, the sum is finite. For $\rho$ small enough $\eta_{\rho}$ satisfies
$\eta_{\rho}\equiv 0\text{ on }\bigcup_{j}\{r_{j}\leq\rho/2\},\qquad\eta_{\rho}\equiv 1\text{ on }B\setminus\bigcup_{j}\{r_{j}\leq\rho\}.$
Define
(54) $\Phi^{(\rho)}:=\eta_{\rho}\,\Phi,\qquad\Phi^{<\rho}:=\Phi-\Phi^{(\rho)},$
and similarly for $\Psi$. Thus $\Phi^{(\rho)},\Psi^{(\rho)}$ vanish in a neighborhood of $D$, while $\Phi^{<\rho},\Psi^{<\rho}$ are supported in a union of wedge neighborhoods.
Definition 5.6 (Renormalized bilinear functional).
For $\tau>0$ and $\Phi,\Psi\in C_{c}^{\infty}(B)$ define
(55) $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau):=\lim_{\rho\downarrow 0}\Bigg[$ $\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$
$\quad+\sum_{j=1}^{N}\int_{U_{j}\times U_{j}}\chi_{j}(b)\chi_{j}(b^{\prime})\,K_{j}^{\mathrm{cone}}(b,b^{\prime};\tau)\,\Phi^{<\rho}(b)\,\Psi^{<\rho}(b^{\prime})\,d\mu_{g^{\mathrm{cone}}_{j}}(b)\,d\mu_{g^{\mathrm{cone}}_{j}}(b^{\prime})\Bigg].$
[…]
$\Phi=\Phi^{(\rho)}+\Phi^{<\rho},\qquad\Psi=\Psi^{(\rho)}+\Psi^{<\rho}$
Figure 3. $\rho$–cutoff decomposition and the gluing mechanism in the definition of the renormalized functional $K_{B}^{\mathrm{ren}}$. The test functions split into an outer part supported in $B_{\mathrm{reg}}$ and an inner part supported near the conic points. The outer contribution uses the base heat kernel $K_{B}$, while the inner contribution is replaced by the model cone kernels $K_{j}^{\mathrm{cone}}$. Taking $\rho\downarrow 0$ yields a canonical cutoff-independent limit (Theorem 5.7).
At this stage one must show that the limit exists and does not depend on any auxiliary choice (wedge charts, cutoffs $\chi_{j}$, or the cutoff profile $\eta$). This is the key intrinsic part of the construction.
Theorem 5.7 (Existence and cutoff-independence).
Assume Assumption 5.1 and the conic parametrix Proposition 5.5. Then for every $\tau>0$ and every $\Phi,\Psi\in C_{c}^{\infty}(B)$, the limit in (55) exists and is finite.
Moreover, $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ is independent of (i) the wedge charts $(r_{j},\theta_{j})$, (ii) the choice of cutoffs $\chi_{j},\widetilde{\chi}_{j}$ used in the parametrix, and (iii) the cutoff profile $\eta$ used to build $\eta_{\rho}$.
Finally, if $\mathrm{supp}\,\Phi\cup\mathrm{supp}\,\Psi\subset B_{\mathrm{reg}}$, then for $\rho$ sufficiently small $\Phi^{<\rho}=\Psi^{<\rho}=0$, hence
$K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)=\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime}).$
Proof.
Fix $0<\rho^{\prime}<\rho<r_{0}$. Subtract the bracketed expressions in (55) at $\rho$ and $\rho^{\prime}$. The difference is supported where at least one variable lies in an annulus $\{\rho^{\prime}/2\leq r_{j}\leq\rho\}$ for some $j$.
On each $U_{j}\times U_{j}$ we insert the decomposition (52). The conic pieces telescope thanks to the way $\Phi^{(\rho)}$ and $\Phi^{<\rho}$ partition $\Phi$ (and similarly for $\Psi$). What remains are integrals of the remainders $R_{j}$ and the measure-comparison errors $d\mu_{g_{B}}-d\mu_{g^{\mathrm{cone}}_{j}}$ against $\Phi,\Psi$ supported in $\{r_{j}<\rho\}$. Using (53) and (49), and integrating against $d\mu_{g^{\mathrm{cone}}_{j}}^{\otimes 2}\sim(r\,dr\,d\theta)(r^{\prime}\,dr^{\prime}\,d\theta^{\prime})$, we obtain
$\Big|\Delta_{\rho,\rho^{\prime}}\Big|\leq C(\tau)\,\|\Phi\|_{L^{\infty}}\|\Psi\|_{L^{\infty}}\int_{0<r,r^{\prime}<\rho}(r+r^{\prime})^{\beta}\,(r\,dr)(r^{\prime}\,dr^{\prime})\ \leq\ C(\tau)\,\rho^{4+\beta}.$
Hence the net is Cauchy as $\rho\downarrow 0$, so the limit exists.
Independence of auxiliary choices follows from the same telescoping estimate: changing charts/cutoffs changes the parametrix only by kernels satisfying the same bound as $R_{j}$, hence the difference between two constructions is $O(\rho^{4+\beta})\to 0$. If supports avoid $D$, then for $\rho$ below the distance to $D$ we have $\Phi^{<\rho}=\Psi^{<\rho}=0$, and (55) collapses to the bare pairing. ∎
5.5. Identification of the total-space bilinear limit with $K_{B}^{\mathrm{ren}}$
We now connect the base renormalized functional to the collapsed total space. Let $(X_{t},g(t))$ be the collapsing family as in Sections 34, with elliptic torus fibration $\Pi_{t}:X_{t}\to B_{\mathrm{reg}}$ on the regular locus, and let $K_{t}(\cdot,\cdot;\tau)$ be the heat kernel on $X_{t}$. We consider the fiber-constant compressed bilinear pairing
(56) $\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}_{t}},\qquad\tau>0,$
where $I_{t}$ is the normalized lift from Section 3 (on the regular locus) and $H_{t}$ is the Friedrichs generator of $E_{t}$.
Assumption 5.8 (Interior collapse input away from $D$).
For every $\rho>0$, define $B(\rho):=B\setminus\bigcup_{j}\{r_{j}\leq\rho\}$, so $B(\rho)\Subset B_{\mathrm{reg}}$. Assume that on $B(\rho)$ the hypotheses of Sections 34 hold uniformly: semi-flat $C^{1}$ control, vertical spectral gap, and hence
(57) $\langle I_{t}\phi,e^{-\tau H_{t}}I_{t}\psi\rangle_{L^{2}(\Pi_{t}^{-1}(B(\rho)))}\longrightarrow\int_{B(\rho)\times B(\rho)}K_{B}(b,b^{\prime};\tau)\,\phi(b)\,\psi(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$
for all $\phi,\psi\in C_{c}^{\infty}(B(\rho))$ and all $\tau>0$.
Theorem 5.9 (Renormalized bilinear limit across $D$).
Assume Assumption 5.8 and the conic hypotheses of this section (Assumption 5.1 and Proposition 5.5). Then for every fixed $\tau>0$ and every $\Phi,\Psi\in C_{c}^{\infty}(B)$,
(58) $\lim_{t\downarrow 0}\mathcal{K}_{t}(\Phi,\Psi;\tau)=K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau).$
Moreover, the convergence is uniform for $\tau$ in compact subsets of $(0,\infty)$.
Proof.
Fix $\tau>0$ and split $\Phi,\Psi$ as in (54) using a cutoff scale $\rho\in(0,r_{0})$:
$\Phi=\Phi^{(\rho)}+\Phi^{<\rho},\qquad\Psi=\Psi^{(\rho)}+\Psi^{<\rho}.$
Expand $\mathcal{K}_{t}(\Phi,\Psi;\tau)$ into four terms.
Step 1: the outer–outer term. Since $\Phi^{(\rho)},\Psi^{(\rho)}$ are supported in $B(\rho/2)\Subset B_{\mathrm{reg}}$, Assumption 5.8 gives
$\lim_{t\downarrow 0}\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{(\rho)};\tau)=\int_{B(\rho/2)\times B(\rho/2)}K_{B}\,\Phi^{(\rho)}\otimes\Psi^{(\rho)}\,d\mu_{g_{B}}^{\otimes 2}.$
Step 2: mixed terms vanish as $\rho\downarrow 0$. Using selfadjointness/contractivity of $e^{-\tau H_{t}}$ and the $L^{2}$-isometry of $I_{t}$,
$\big|\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{<\rho};\tau)\big|\leq\|I_{t}\Phi^{(\rho)}\|_{L^{2}_{t}}\,\|I_{t}\Psi^{<\rho}\|_{L^{2}_{t}}=\|\Phi^{(\rho)}\|_{L^{2}(B)}\,\|\Psi^{<\rho}\|_{L^{2}(B)}.$
Since $\Psi^{<\rho}$ is supported in $\bigcup_{j}\{r_{j}<\rho\}$ and $\Psi$ is smooth with compact support, we have $\|\Psi^{<\rho}\|_{L^{2}(B)}\to 0$ as $\rho\downarrow 0$, hence the mixed term tends to 0 uniformly in $t$. The same estimate applies to $\mathcal{K}_{t}(\Phi^{<\rho},\Psi^{(\rho)};\tau)$.
Step 3: the inner–inner term is captured by the cone model up to an error vanishing with $\rho$. Insert a partition of unity $\sum_{j}\chi_{j}$ on $\mathrm{supp}(\Phi^{<\rho})\cup\mathrm{supp}(\Psi^{<\rho})$ and work wedgewise. On each $U_{j}$ compare the base kernel $K_{B}$ with the model cone kernel via Proposition 5.5. The remainder estimate (53) and the measure comparison (49) show that the error made by replacing $K_{B}$ with $\chi_{j}\otimes\chi_{j}\,K_{j}^{\mathrm{cone}}$ in the inner–inner pairing is $O_{\tau}(\rho^{4+\beta})$. Combining with interior convergence on truncated sets $B(\rho^{\prime})$ and then letting $\rho^{\prime}\downarrow 0$ yields
$\lim_{t\downarrow 0}\mathcal{K}_{t}(\Phi^{<\rho},\Psi^{<\rho};\tau)=\sum_{j=1}^{N}\int_{U_{j}\times U_{j}}\chi_{j}\otimes\chi_{j}\,K_{j}^{\mathrm{cone}}\,\Phi^{<\rho}\otimes\Psi^{<\rho}\,d\mu_{g^{\mathrm{cone}}_{j}}^{\otimes 2}\;+\;O_{\tau}(\rho^{4+\beta}).$
Step 4: send $\rho\downarrow 0$. Combining Steps 1–3 and using the definition (55) plus Theorem 5.7 yields (58). ∎
5.6. Section summary
We have established the following: (i) wedge charts and $C^{1}$ conic control of $g_{B}$ near $D$; (ii) a conic parametrix for $K_{B}$ with $(r+r^{\prime})^{\beta}$-weighted remainder bounds; (iii) the renormalized bilinear functional $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ together with its cutoff-independence; and (iv) the identification of the collapsed total-space bilinear pairing limit with $K_{B}^{\mathrm{ren}}$ for each fixed $\tau>0$.
In the next section we will streamline the hypotheses needed in Assumption 5.8 in the geometric semi-flat collapse setting, and record quantitative error bounds (in $t$) for the renormalized bilinear convergence.
6. Quantitative estimates and error bookkeeping
In this section we combine three mechanisms into a single quantitative comparison: (i) the interior (semi-flat) collapse estimate with rate $\epsilon_{t}\sim e^{-c/s^{*}(t)}$; (ii) the edge conic parametrix remainder of size $O(\rho^{4+\beta})$; and (iii) the mixed decoupling controlled by the $L^{2}$-smallness of the inner cutoffs (with off-diagonal bounds such as Davies–Gaffney; see [7, 16]). We then state a final citable theorem encoding the iterated limit “first $t\downarrow 0$, then $\rho\downarrow 0$”.
6.1. Two-parameter truncation and the $\rho$–renormalized target
Retain the wedge charts $(U_{j},r_{j},\theta_{j})$, the cutoffs $\chi_{j}$, and the cutoff decomposition $\Phi=\Phi^{(\rho)}+\Phi^{<\rho}$ from §5. Fix $\tau>0$ and $\rho\in(0,r_{0})$.
Total-space bilinear pairing.
Let $(X_{t},g(t))$ be the collapsing family and let $H_{t}$ be the Friedrichs generator of the Dirichlet form $E_{t}$ (with the boundary convention fixed in §3–§4). For $\Phi,\Psi\in C_{c}^{\infty}(B)$ define the compressed (fiber-constant) pairing
(59) $\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,\,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}(X_{t})},$
where $I_{t}$ is the normalized lift.
$\rho$–renormalized base target.
Define the $\rho$–renormalized target functional
(60) $K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau):=$ $\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$
$\quad+\sum_{j=1}^{N}\int_{U_{j}\times U_{j}}\chi_{j}(b)\chi_{j}(b^{\prime})\,K^{\mathrm{cone}}_{j}(b,b^{\prime};\tau)\,\Phi^{<\rho}(b)\,\Psi^{<\rho}(b^{\prime})\,d\mu_{g^{\mathrm{cone}}_{j}}(b)\,d\mu_{g^{\mathrm{cone}}_{j}}(b^{\prime}).$
By Theorem 5.7, $K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)$ converges as $\rho\downarrow 0$ to $K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)$, independently of auxiliary choices.
6.2. Quantitative interior estimate at fixed $\rho$
Let $\epsilon_{t}:=Ce^{-c/s^{*}(t)}$ be the semi-flat error gauge from §3–§4. Fix $\rho>0$ and consider the buffer region
$B(\rho/2):=B\setminus\bigcup_{j=1}^{N}\{r_{j}\leq\rho/2\}\Subset B_{\mathrm{reg}}.$
Then $\Phi^{(\rho)},\Psi^{(\rho)}\in C_{c}^{\infty}(B(\rho/2))$, and all interior hypotheses hold uniformly on $B(\rho/2)$.
Proposition 6.1 (Interior comparison with exponential rate).
Fix $\rho\in(0,r_{0})$ and a compact time window $\tau\in[\tau_{1},\tau_{2}]\subset(0,\infty)$. There exists $C=C(\rho,\tau_{1},\tau_{2})$ such that for all sufficiently small $t$,
(61) $\Big|\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{(\rho)};\tau)-\!\!\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}\!\!K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}^{\otimes 2}\Big|\leq C\,\epsilon_{t}\,\|\Phi\|_{L^{2}(B)}\|\Psi\|_{L^{2}(B)}.$
The estimate holds uniformly for $\tau\in[\tau_{1},\tau_{2}]$.
Proof.
On the fixed precompact set $B(\rho/2)$ we have the quantitative compressed semigroup bound
$\|P_{t}e^{-\tau H_{t}}I_{t}-e^{-\tau H_{B}}\|_{L^{2}\to L^{2}}\leq C(\rho,\tau_{1},\tau_{2})\,\epsilon_{t},\qquad\tau\in[\tau_{1},\tau_{2}],$
from §4 (under the exponentially small semi-flat error hypothesis). Pairing against $\Phi^{(\rho)},\Psi^{(\rho)}$ and using the $L^{2}$ control for $I_{t}$ yields (61). ∎
6.3. Mixed-term control
We bound the mixed pairings by contractivity and the $L^{2}$-smallness of the inner cutoffs.
Lemma 6.2 (Mixed terms: contractive bound).
For every $\tau>0$, all sufficiently small $t$, all $\rho\in(0,r_{0})$, and all $\Phi,\Psi\in C_{c}^{\infty}(B)$,
(62) $\Big|\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{<\rho};\tau)\Big|$ $\leq C\,\|\Phi^{(\rho)}\|_{L^{2}(B)}\,\|\Psi^{<\rho}\|_{L^{2}(B)},$
(63) $\Big|\mathcal{K}_{t}(\Phi^{<\rho},\Psi^{(\rho)};\tau)\Big|$ $\leq C\,\|\Phi^{<\rho}\|_{L^{2}(B)}\,\|\Psi^{(\rho)}\|_{L^{2}(B)}.$
Here $C>0$ is independent of $t,\rho$ (for $t$ small).
Proof.
By $L^{2}$-contractivity of $e^{-\tau H_{t}}$ and Cauchy–Schwarz,
$\big|\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{<\rho};\tau)\big|\leq\|I_{t}\Phi^{(\rho)}\|_{L^{2}(X_{t})}\,\|I_{t}\Psi^{<\rho}\|_{L^{2}(X_{t})}.$
Using the $L^{2}$ control of $I_{t}$ (uniformly for $t$ small) yields (62). The second estimate is identical. ∎
Since in dimension two one has $\|\Phi^{<\rho}\|_{L^{2}(B)}=O(\rho)\|\Phi\|_{L^{\infty}(B)}$, the mixed terms vanish as $\rho\downarrow 0$.
Lemma 6.3 ($L^{2}$ size of the inner cutoff).
There exists $C>0$ such that for all sufficiently small $\rho$,
(64) $\|\Phi^{<\rho}\|_{L^{2}(B)}\leq C\,\rho\,\|\Phi\|_{L^{\infty}(B)},\qquad\|\Psi^{<\rho}\|_{L^{2}(B)}\leq C\,\rho\,\|\Psi\|_{L^{\infty}(B)}.$
Proof.
$\Phi^{<\rho}$ is supported in $\bigcup_{j}\{r_{j}<\rho\}$, whose $g_{B}$–area is $O(\rho^{2})$ by the conic volume form $d\mu_{g_{B}}=(1+O(r^{\beta}))\,\alpha_{j}r\,dr\,d\theta$ near each $p_{j}$. Hence
$\|\Phi^{<\rho}\|_{L^{2}(B)}^{2}\leq\|\Phi\|_{L^{\infty}(B)}^{2}\,\mu_{g_{B}}\Big(\bigcup_{j}\{r_{j}<\rho\}\Big)\leq C\,\rho^{2}\,\|\Phi\|_{L^{\infty}(B)}^{2},$
and similarly for $\Psi^{<\rho}$. ∎
6.4. Edge parametrix remainder: the $O(\rho^{4+\beta})$ term
We quantify the convergence of $K^{\mathrm{ren}}_{B,\rho}$ to $K^{\mathrm{ren}}_{B}$ and isolate the $\rho^{4+\beta}$ term coming from the weighted remainder in the conic parametrix.
Proposition 6.4 (Edge remainder bound).
Assume the conic asymptotics of §5 with exponent $\beta>0$. Fix a compact time window $\tau\in[\tau_{1},\tau_{2}]\subset(0,\infty)$. Then there exists $C=C(\tau_{1},\tau_{2})$ such that for all $\rho\in(0,r_{0})$ and all $\Phi,\Psi\in C_{c}^{\infty}(B)$,
(65) $\big|K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\big|\leq C\,\rho^{4+\beta}\,\|\Phi\|_{L^{\infty}(B)}\,\|\Psi\|_{L^{\infty}(B)},\qquad\tau\in[\tau_{1},\tau_{2}].$
Proof.
This is the quantitative version of the Cauchy estimate in Theorem 5.7. Using the remainder bound (53) and the measure comparison near each $p_{j}$, the difference between two cutoff radii is controlled by
$C(\tau_{1},\tau_{2})\,\|\Phi\|_{\infty}\|\Psi\|_{\infty}\int_{0<r,r^{\prime}<\rho}(r+r^{\prime})^{\beta}(r\,dr)(r^{\prime}\,dr^{\prime})\ \lesssim\ \rho^{4+\beta}\|\Phi\|_{\infty}\|\Psi\|_{\infty},$
uniformly for $\tau\in[\tau_{1},\tau_{2}]$. ∎
6.5. Final bookkeeping and the citable iterated-limit theorem
We can now combine the interior estimate, the mixed-term control, and the edge remainder into one statement.
Theorem 6.5 (Quantitative renormalized comparison and iterated limit).
Assume:
  1. (1)
    (Semi-flat interior with exponential rate) On every compact subset of $B_{\mathrm{reg}}$ the semi-flat collapse hypotheses of §3–§4 hold with error $\epsilon_{t}\leq Ce^{-c/s^{*}(t)}$ and yield the quantitative compressed semigroup estimate used in Proposition 6.1.
  2. (2)
    (Conic base model) The base metric satisfies the $C^{1}$ conic asymptotics of Assumption 5.1 with exponent $\beta>0$, and the conic parametrix Proposition 5.5 holds on a time interval containing $\tau_{1},\tau_{2}]$.
Fix a compact time window $\tau\in[\tau_{1},\tau_{2}]\subset(0,\infty)$. Then there exist constants $C>0$ such that for all sufficiently small $t$ and all $\rho\in(0,r_{0})$,
(66) $\Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\Big|\leq\;$ $C(\rho,\tau_{1},\tau_{2})\,\epsilon_{t}\,\|\Phi\|_{L^{2}(B)}\|\Psi\|_{L^{2}(B)}$
$\quad+C\,\Big(\|\Phi^{(\rho)}\|_{L^{2}(B)}\|\Psi^{<\rho}\|_{L^{2}(B)}+\|\Phi^{<\rho}\|_{L^{2}(B)}\|\Psi^{(\rho)}\|_{L^{2}(B)}\Big)$
$\quad+C\,\rho^{4+\beta}\,\|\Phi\|_{L^{\infty}(B)}\|\Psi\|_{L^{\infty}(B)}.$
Moreover, since $K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\to K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)$ as $\rho\downarrow 0$, we have the iterated-limit identity
(67) $\lim_{\rho\downarrow 0}\ \limsup_{t\downarrow 0}\ \Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\Big|=0,\qquad\tau\in[\tau_{1},\tau_{2}].$
In particular, $\mathcal{K}_{t}(\Phi,\Psi;\tau)\to K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)$ for each fixed $\tau>0$.
Proof.
Expand $\Phi=\Phi^{(\rho)}+\Phi^{<\rho}$ and $\Psi=\Psi^{(\rho)}+\Psi^{<\rho}$ in (59). This yields four terms. The outer–outer term is compared to the base $K_{B}$ pairing with rate $\epsilon_{t}$ by Proposition 6.1. The mixed terms are bounded by Lemma 6.2. For the inner–inner contribution, we add and subtract the conic model pairing; the difference is controlled by the conic parametrix remainder estimate, yielding the $C\rho^{4+\beta}$ term (as in Proposition 6.4). Collecting these bounds gives (66).
To obtain (67), fix $\rho>0$ and let $t\downarrow 0$: the first term vanishes since $\epsilon_{t}\to 0$. Then let $\rho\downarrow 0$: by Lemma 6.3 the mixed term is $O(\rho)$, and the edge term is $O(\rho^{4+\beta})$. Finally use Proposition 6.4 to replace $K^{\mathrm{ren}}_{B,\rho}$ with $K^{\mathrm{ren}}_{B}$. ∎
Remark 6.6 (A one-parameter corollary (optional)).
If one wishes to turn the two-parameter bookkeeping into a single $t\downarrow 0$ estimate, one may choose a function $\rho=\rho(t)\downarrow 0$ and combine (66), Lemma 6.3 and Proposition 6.4. For instance, choosing $\rho(t):=\epsilon_{t}^{1/(4+\beta)}$ gives a concrete (non-optimal) bound
$\Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\Big|\leq C\Big(\epsilon_{t}+\epsilon_{t}^{1/(4+\beta)}\Big),$
for fixed $\tau\in[\tau_{1},\tau_{2}]$ and bounded $\Phi,\Psi$. We keep the iterated-limit formulation (67) as the canonical statement.
Remark 6.7 (On small-time information).
Theorem 6.5 is a fixed-$\tau$ statement and is designed to justify the renormalized limit across $D$. Any additional analysis of the behavior of $K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)$ as $\tau\downarrow 0$ is purely base/model-driven and is not used in the convergence arguments above.
7. Synthesis: renormalized limit for compressed heat flow
We now assemble the two complementary outputs established in Theorem 5.7 (cutoff-independence and canonicity of $K_{B}^{\mathrm{ren}}$) and Theorem 6.5 (quantitative comparison and the fixed-$\tau$ collapsed limit) into a single citable statement.
Theorem 7.1 (Main theorem: renormalized compressed heat kernel limit).
Assume:
  1. (1)
    (Semi-flat collapse and vertical gap on $B_{\mathrm{reg}}$) on every precompact $B_{0}\Subset B_{\mathrm{reg}}$ the family $(X_{t},g(t))$ admits semi-flat control (with $\varepsilon_{t}\to 0$; in addition one may assume an exponential gauge $\epsilon_{t}\leq Ce^{-c/s^{*}(t)}$ when stating explicit rates), and the fiberwise first nonzero eigenvalue satisfies $\lambda_{1}(F_{b,t})\gtrsim s^{*}(t)^{-2}$.
  2. (2)
    (Conic base structure at the discriminant) the base metric $g_{B}$ satisfies the $C^{1}$ conic asymptotics near each $p_{j}\in D$ as in Assumption 5.1, so that the cutoff-independent functional $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ of Definition 5.6 is well-defined.
  3. (3)
    (Off-diagonal control and bookkeeping framework) the Davies–Gaffney decoupling and the quantitative bookkeeping hypotheses used in Section 6 hold on the relevant truncations away from $D$.
Fix a boundary type among Dirichlet/Neumann (and Robin with nonnegative weight) as in Sections 46, and let $H_{t}$ denote the corresponding Friedrichs generator on the truncated domains (with normalized lift $I_{t}$).
(i) Renormalized bilinear limit at fixed time. For every $\tau>0$ and every $\Phi,\Psi\in C_{c}^{\infty}(B)$,
(68) $\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,\,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}(X_{t})}\longrightarrow K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)\qquad(t\downarrow 0).$
If $\operatorname{supp}\Phi\cup\operatorname{supp}\Psi\subset B_{\mathrm{reg}}$, then the renormalization is trivial and $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)$ coincides with the usual base pairing against $K_{B}(\cdot,\cdot;\tau)$.
(ii) Quantitative bookkeeping (two-parameter form). Fix a compact time window $\tau\in[\tau_{1},\tau_{2}]\subset(0,\tau_{0}]$ as in Section 6. Then for all sufficiently small $t$ and all $\rho\in(0,r_{0})$ one has
(69) $\big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\big|\leq\mathrm{Err}_{\mathrm{int}}(t;\Phi,\Psi)+\mathrm{Err}_{\mathrm{mix}}(\rho,\tau;\Phi,\Psi)+\mathrm{Err}_{\mathrm{edge}}(\rho,\tau;\Phi,\Psi),$
with the same error structure as in Theorem 6.5. Consequently,
$\lim_{\rho\downarrow 0}\limsup_{t\downarrow 0}\big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)\big|=0\quad\text{uniformly for }\tau\in[\tau_{1},\tau_{2}].$
Proof.
This is exactly Theorem 6.5, together with the canonicity of $K_{B}^{\mathrm{ren}}$ from Theorem 5.7. ∎
Corollary 7.2 (Kernel pairing formulation).
On each precompact truncation where the heat kernel representation holds (Section 4), the convergence (68) is equivalent to the renormalized bilinear heat-kernel pairing limit
$\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{g(t)}(x)\,d\mu_{g(t)}(y)\longrightarrow K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau),\qquad t\downarrow 0,$
with the same error structure as in (69) after inserting the $\rho$–decomposition.
8. Geometric input in semi-flat elliptic fibrations
This section records a concrete and verifiable set of geometric conditions implying the analytic hypotheses used in Sections 37. The goal is not to reprove semi-flat geometry, but to isolate the minimal inputs needed to apply Theorem 7.1 in geometric collapse problems.
8.1. Semi-flat approximation and the identification operators
Let $\Pi_{t}:X_{t}\to B$ be a smooth elliptic fibration over $B_{\mathrm{reg}}:=B\setminus D$ with flat torus fibers $F_{b,t}\cong\mathbb{T}^{2}$ and a family of Riemannian metrics $g(t)$ on $X_{t}$. Fix a precompact domain $B_{0}\Subset B_{\mathrm{reg}}$ and set $X_{t,0}:=\Pi_{t}^{-1}(B_{0})$.
Assume that on $X_{t,0}$ there exists a product reference metric
$g_{t}^{\Pi}:=g_{B}\oplus g_{F,t},$
as in the semi-flat collapse regimes studied in [2, 10, 11], where $g_{B}$ is a fixed smooth metric on $B_{0}$ and $g_{F,t}$ restricts to a flat metric on each fiber, such that
(70) $\|g(t)-g_{t}^{\Pi}\|_{C^{1}(X_{t,0},\,g_{t}^{\Pi})}\leq\epsilon_{t},\qquad\epsilon_{t}\downarrow 0\quad(\text{and optionally }\epsilon_{t}\leq Ce^{-c/s^{*}(t)}).$
Let $A_{t}(b):=\mu_{F_{b,t}}(F_{b,t})$ be the fiber area and let $d\vartheta_{b,t}$ denote the Haar probability measure on $(F_{b,t},g_{F,t})$. Quantitative disintegration of volume along $\Pi_{t}$ yields measurable $\rho_{t}>0$ with fiberwise normalization $\int_{F_{b,t}}\rho_{t}\,d\vartheta_{b,t}=1$ such that for all $f\in L^{1}(X_{t,0})$,
$\int_{X_{t,0}}f\,d\mu_{g(t)}=\int_{B_{0}}A_{t}(b)\left(\int_{F_{b,t}}f\,\rho_{t}\,d\vartheta_{b,t}\right)d\mu_{g_{B}}(b).$
Moreover, the $C^{1}$ semi-flat control (70) implies the quantitative bounds
$\|\rho_{t}-1\|_{L^{\infty}(X_{t,0})}+\|\nabla_{g_{B}}\rho_{t}\|_{L^{\infty}(X_{t,0})}+\|\nabla_{g_{B}}\log A_{t}\|_{L^{\infty}(B_{0})}\ \leq\ C\,\epsilon_{t},$
after choosing the standard horizontal/vertical splitting associated with $g_{t}^{\Pi}$.
The normalized lift/average operators are then
$(I_{t}v)(x)=A_{t}(\Pi_{t}x)^{-1/2}v(\Pi_{t}x),\qquad(P_{t}u)(b)=A_{t}(b)^{1/2}\int_{F_{b,t}}u\,\rho_{t}\,d\vartheta_{b,t},$
so that $P_{t}I_{t}=\mathrm{Id}$ and $I_{t}=P_{t}^{*}$, with two-sided $L^{2}$ near-isometry and $H^{1}$ compatibility on $B_{0}$. These are exactly the structural inputs used to formulate Mosco convergence and compressed semigroup limits.
8.2. Vertical spectral gap from fiber scale
The collapse scale $s^{*}(t)\downarrow 0$ is encoded in a uniform upper bound on the fiber diameter:
(71) $\mathrm{diam}(F_{b,t},g_{F,t})\leq s^{*}(t)\qquad\text{for all }b\in B_{0}.$
Lemma 8.1 (Flat torus gap from diameter).
There exists a universal constant $c_{0}>0$ such that for every flat 22–torus $(\mathbb{T}^{2},h)$,
$\lambda_{1}(\mathbb{T}^{2},h)\geq c_{0}\,\mathrm{diam}(\mathbb{T}^{2},h)^{-2},$
where $\lambda_{1}$ denotes the first nonzero eigenvalue of $\Delta_{h}$. Consequently, under (71),
(72) $\lambda_{1}(F_{b,t},g_{F,t})\geq c_{0}\,s^{*}(t)^{-2}\qquad\text{uniformly for }b\in B_{0}.$
Proof.
This follows, for instance, from the Zhong–Yang lower bound $\lambda_{1}\geq\pi^{2}/\mathrm{diam}^{2}$ on compact manifolds with $\mathrm{Ric}\geq 0$ [23], applied to flat tori (for which $\mathrm{Ric}=0$). ∎
Lemma 8.1 yields the fiberwise Poincaré inequality and the suppression of fiber-nonconstant modes at energies $\ll s^{*}(t)^{-2}$, which is the starting point of the analytic chain in Sections 34.
8.3. Conic base input at the discriminant
Near each discriminant point $p_{j}\in D$, fix a wedge chart $U_{j}$ with polar coordinate $r=r_{j}$ and assume that $g_{B}$ is $C^{1}$–close to the flat cone metric
$g_{j}^{\mathrm{cone}}:=dr^{2}+\alpha_{j}^{2}r^{2}\,d\theta^{2}$
in the sense of Assumption 5.1, with some exponent $\beta>0$ controlling the coefficient defect. This is precisely the hypothesis needed to construct the conic parametrix and the cutoff-independent functional $K_{B}^{\mathrm{ren}}(\cdot,\cdot;\tau)$ (Section 5).
8.4. Application template
In summary, to apply Theorem 7.1 it suffices to verify on each $B_{0}\Subset B_{\mathrm{reg}}$:
  1. (1)
    semi-flat $C^{1}$ control (70) for $g(t)$ relative to $g_{B}\oplus g_{F,t}$ (with $\epsilon_{t}\to 0$);
  2. (2)
    the induced disintegration bounds on $(A_{t},\rho_{t})$ including $\|\nabla_{g_{B}}\rho_{t}\|_{\infty}=O(\epsilon_{t})$;
  3. (3)
    fiber scale control (71), hence the vertical gap (72);
  4. (4)
    a $C^{1}$ conic model for $g_{B}$ near each $p_{j}\in D$ with cone parameter $\alpha_{j}$.
Once these are in place, the analytic mechanism (vertical suppression $\Rightarrow$ Mosco $\Rightarrow$ compressed semigroups $\Rightarrow$ interior bilinear kernel limit $\Rightarrow$ conic renormalization) applies verbatim, with the quantitative bookkeeping recorded in Section 6.